### v1

# Delays in Spiking Neural Networks: A State Space Model Approach

Delays in Spiking Neural Networks:
A State Space Model Approach
Sanja Karilanova, Subhrakanti Dey, Ayça Özçelikkale
Department of Electrical Engineering, Uppsala University, Sweden
{Sanja.Karilanova, Subhrakanti.Dey, Ayca.Ozcelikkale}@angstrom.uu.se
S. Karilanova and A. Özçelikkale acknowledge the support of Center for Interdisciplinary Mathematics (CIM), and AI4Research, Uppsala University.The computations were enabled by resources provided by the National Academic Infrastructure for Supercomputing in Sweden (NAISS), partially funded by the Swedish Research Council through grant agreement no. 2022-06725.
I Abstract
Spiking neural networks (SNNs) are biologically inspired, event-driven models that are suitable for processing temporal data and offer energy-efficient computation when implemented on neuromorphic hardware. In SNNs, richer neuronal dynamic allows capturing more complex temporal dependencies, with delays playing a crucial role by allowing past inputs to directly influence present spiking behavior. We propose a general framework for incorporating delays into SNNs through additional state variables. The proposed mechanism enables each neuron to access a finite temporal input history. The framework is agnostic to neuron models and hence can be seamlessly integrated into standard spiking neuron models such as LIF and adLIF. We analyze how the duration of the delays and the learnable parameters associated with them affect the performance. We investigate the trade-offs in the network architecture due to additional state variables introduced by the delay mechanism. Experiments on the Spiking Heidelberg Digits (SHD) dataset show that the proposed mechanism matches the performance of existing delay-based SNNs while remaining computationally efficient. Moreover, the results illustrate that the incorporation of delays may substantially improve performance in smaller networks. Code available at https://github.com/sannkka/Modeling-State-Delays-in-SNNs.
II Introduction
Spiking Neural Networks (SNNs) [MAASS19971659] have emerged as a more biologically plausible class of neural models compared to conventional artificial neural networks (ANNs), offering promising advantages in temporal processing and energy efficiency, particularly when deployed on neuromorphic hardware [davies_advancing_2021]. SNNs encode information in the timing of discrete spikes, allowing them to capture temporal structure and making them well suited for modeling time-dependent signals such as audio [bittar2022surrogate], sensory [10.3389/fnins.2022.951164], or visual patterns [Gallego2022]. Despite these advantages, achieving performance comparable to that of ANNs remains an open challenge, in part due to the limited temporal expressivity of current SNN models.
In biological neural systems, temporal delays play a critical computational role. Delays that enable neurons to integrate information from multiple past moments support tasks such as hearing [carr1993processing] and brain activity [carr1990circuit]. Inspired by these findings, numerous works have explored incorporating delays into computational SNNs models. Theoretical analyses have shown that an SNN with $k$ adjustable delays can compute a much richer class of functions than a threshold circuit with $k$ adjustable weights [MAASS199926]. Empirical studies have also demonstrated improved performance when optimizing spike transmission delays along with synaptic weights or membrane time constants [deckers2024co, sun2025exploitingheterogeneousdelaysefficient]. These results suggest that delays can enrich the information processing capabilities of SNNs and their internal representation. However, despite these promising results, practical methods for introducing structured, scalable, and differentiable forms of delay into modern deep SNNs remains an ongoing research problem.
In this work, we propose a method based on State Space Models (SSMs) for equipping SNNs with delays. The method uses a time-shift based state transition, which enables each neuron to access both its current input and the most recent history of past input values.The formulation is general and can be incorporated into arbitrary spiking neuron models. Conceptually, the framework bridges ideas from SSMs and SNNs, allowing SNNs to capture a fixed number of past inputs in a structured and interpretable manner.
Our contributions can be summarized as follows:
  • We propose a method for incorporating delays within the state dynamics through additional delay variables, enabling neurons to access a finite temporal history of inputs.
  • The proposed formulation is general and easily integrable into existing spiking neuron models, including LIF, adLIF, and RadLIF.
  • We conduct a systematic analysis of delay order, i.e. the number of allowed past inputs, and their associated learnable parameters, and study the trade-offs between the number of neurons in the network and the delay order.
  • The proposed framework introduces either no additional trainable parameters (i.e when the delay parameters are not trained) or a number of trainable parameters comparable to the other SNN delay models.
The proposed models achieve performance comparable to existing delay-based SNN approaches on the Spiking Heidelberg Digits (SHD) dataset. Moreover, the observed trade-offs highlight promising directions for future exploration, particularly showing that incorporating delays can substantially improve accuracy in smaller SNNs where the neuron count limits temporal capacity.
Figure 1: Illustration of different approaches of modeling delays. Our proposed delays are modeled within the state dynamics of the neuron (green). Figure is inspired by [DelGrad, Fig. 3(a)].
III Related Work
In the SNN literature, three main types of delays are commonly modeled: axonal delays, applied to the output of a neuron [shrestha2018slayer, 10.3389/fnins.2023.1275944, sun2025exploitingheterogeneousdelaysefficient]; dendritic delays, applied to its input [diaz2016efficient]; and synaptic delays, specific to each connection between neuron pairs [10.1145/3589737.3606009, deckers2024co, hammouamri2023learning, dominijanni2025extendingspiketimingdependentplasticity, meszaros2024learning]. In contrast to these formulations, we propose an SSM based approach where delays are modeled within the state dynamics of the neuron. See Figure 1 for a schematic illustration.
The most common ways to model delays in SNNs is to learn a single delay value applied to each spike [shrestha2018slayer], or to employ temporal convolutions [hammouamri2023learning]. Several works explore the role of delays in SNNs by jointly training weights and delays [deckers2024co], learning only delays [10.1145/3589737.3606009], or showing that heterogeneous delays can reduce weight-precision requirements [sun2025exploitingheterogeneousdelaysefficient]. Others relate delays to temporal parameters such as time constants [10.1371/journal.pcbi.1012673] or study their role in the formation of temporal hierarchies between layers [moro2024roletemporalhierarchyspiking]. In these works, various learning algorithms have been used, such as variations of backpropagation through time (BPTT) [hammouamri2023learning, shrestha2018slayer, queant2025delreclearningdelaysrecurrent, meszaros2024learning], biologically inspired spike-timing-dependent plasticity (STDP) [dominijanni2025extendingspiketimingdependentplasticity], and exact spike-time gradients [DelGrad, meszaros2025efficient]. Our formulation represents delays as state variables evolving over time. This structure remains compatible with BPTT, enabling training with the popular surrogate-gradient methods [neftci2019surrogate] used to train SNNs.
Delays have long been studied in fields such as signal processing and control theory. In discrete-time dynamical systems, they are typically modeled through delay-difference or state-space formulations, where past states are explicitly included to capture finite memory effects [RICHARD20031667, SenameOlivier2005Sots]. Recently, modern deep SSMs, which stack multiple state-space layers combined with nonlinearities, have demonstrated strong performance in sequence modeling tasks [gu2022efficientlymodelinglongsequences, Gupta_DSS]. Even more recently, parameterization has been introduced in deep SSMs, effectively capturing delayed state dependencies [fu2022hungry, zhang2023effectively].
We build on this formulation to represent controllable delays within spiking neural network models. A common approach in modern SNNs embeds spiking dynamics within SSMs, where at each time step the output of the deep SSM layer is passed through a firing threshold function [stan2023learning, bal2024rethinkingspikingneuralnetworks, karilanova2025statespacemodelinspiredmultipleinput, karilanova_reset, du2024spiking]. We continue in this direction by utilizing the SSM perspective on delay modeling to spiking neural architectures, enabling SNNs to incorporate structured temporal dependencies through the same principled state-space formulation.
IV Methods
IV-A Preliminaries
IV-A1 General Spiking Neuron Model
A single general feed-forward discrete-time $n$-state dimensional spiking neuron can be defined as [karilanova2025zeroshottemporalresolutiondomain, karilanova2025statespacemodelinspiredmultipleinput, karilanova_reset]:
$\!\bm{v}_{s}[t+1]$ $=\bm{A}_{s}\bm{v}_{s}[t]+\bm{B}_{s}i_{s}[t]-\bm{R}s[t]$ (1a)
$s[t]$ $=f_{\bf{\Theta}}(\bm{v}_{s}[t])=\begin{cases}1\text{ if }\bm{v}_{s}[t]\in\bf{\Theta}\\ 0\text{ otherwise}\end{cases}$ (1b)
where $i_{s}[t]\in\mathbb{R}^{1\times 1}$ is the input to the neuron, $s[t]\in\mathbb{R}^{1\times 1}$ is the output of the neuron and $\bm{v}_{s}\in\mathbb{R}^{n_{s}\times 1}$ is the state variable of the neuron. The matrices $\bm{A}_{s}\in\mathbb{R}^{n_{s}\times n},\bm{B}_{s}\in\mathbb{R}^{n_{s}\times 1},\bm{R}\in\mathbb{R}^{n_{s}\times 1}$ represents the leak, impact of the input and the reset feedback, respectively. The spiking behavior is described by $f_{\bf{\Theta}}(\cdot)$, where the neuron spikes when the state enters into the region described by $\bf{\Theta}$.
The neuron model in (1) generalizes several commonly used formulations, including the Leaky Integrate-and-Fire (LIF) and adaptive LIF (adLIF) neurons, see [karilanova2025zeroshottemporalresolutiondomain, karilanova2025statespacemodelinspiredmultipleinput, karilanova_reset] for illustration of such representations.
Spiking neural networks are built by connecting neurons such that the output spikes from the previous (pre-synaptic, l1l-1) layer serve as inputs to the next (post-synaptic, $l$) layer. Let $L_{\text{pre}}$ and $L_{\text{post}}$ be the number of neurons in the pre and post synaptic layer. Let $\bm{W}\in\mathbb{R}^{L_{\text{post}}\times L_{\text{pre}}}$ be the weight matrix connecting the two layers, and $\bm{V}\in\mathbb{R}^{L_{\text{post}}\times L_{\text{post}}}$ be the weight matrix associated with the recurrent connections between neurons in a layer. Then $\bm{i}[t]=\bm{W}\times\bm{s}^{l-1}[t]+\bm{V}\times\bm{s}^{l}[t]$, where $\bm{s}^{l-1}[t]$ and $\bm{s}^{l}[t]$ are a vector of all the spikes from the previous and current layer, respectively.
IV-A2 State Space Model with Time Shift State Transition Matrix
A discrete-time time-invariant linear SSM can be written as [ljung_SI]:
$\bm{v}_{d}[t+1]=\bm{A}_{d}\bm{v}_{d}[t]+\bm{B}_{d}i_{d}[t]$ (2a)
$\bm{y}_{d}[t]=\bm{C}_{d}\bm{v}_{d}[t]+\bm{D}_{d}i_{d}[t],$ (2b)
where $\bm{v}_{d}[t],i_{d}[t],\bm{y}_{d}[t]$ denote the state vector, the input vector, and the output vector, respectively. Here, the state transition matrix $\bm{A}_{d}\in\mathbb{R}^{n_{d}\times n_{d}}$ describes the internal recurrent behavior of SSM. The matrices $\bm{B}_{d}\in\mathbb{R}^{n_{d}\times 1},\bm{C}_{d}\in\mathbb{R}^{n_{out}\times n_{d}},\bm{D}_{d}\in\mathbb{R}^{n_{out}\times 1}$ describe the interaction of the SSM through its inputs and outputs.
The lower shift matrix is a structured form of the state-transition matrix with nonzero elements only on the subdiagonal. Let $\bm{A}_{d}$ be the lower shift matrix:
$\bm{A}_{d}$ $=\begin{bmatrix}0&0&\ldots&0&0\\ a_{1}&0&\ldots&0&0\\ 0&a_{2}&\ldots&0&0\\ \vdots&&\ddots&\vdots&\vdots\\ 0&0&\ldots&a_{n_{d}-1}&0\end{bmatrix}\text{i.e.,}$ (3a)
$[\bm{A}_{d}]_{i,j}$ $=\begin{cases}a_{j}&\text{ for }i-1=j\\ 0&\text{ otherwise }\end{cases}$ (3b)
where $\bm{A}_{d}]_{i,j}$ represents the $i$th row $j$th column element of the matrix $\bm{A}_{d}$. Let
$\bm{B}_{d}=[1,0,\ldots,0]^{T}.$ (4)
Then for the SSM state variable $\bm{v}_{d}$ defined in (2) and initial state $\bm{v}_{d}[0]=\bm{0}^{n_{d}\times 1}$ we have:
$\bm{v}_{d}[1]$ $=[i_{d}[1],\,0,\,0,\,...,\,0]^{T}$
$\bm{v}_{d}[2]$ $=[i_{d}[2],\,a_{1}i_{d}[1],\,0,\,...,\,0]^{T}$
$\bm{v}_{d}[3]$ $=[i_{d}[3],\,a_{1}i_{d}[2],\,a_{2}a_{1}i_{d}[1],\,...,\,0]^{T}$
$\,\,\,\vdots$
$\bm{v}_{d}[t]$ $=[i_{d}[t],\,a_{1}i_{d}[t-1],\,a_{2}a_{1}i_{d}[t-2],\,$...,\,
$\hskip 85.35826pt(\prod^{n-1}_{k=1}a_{k})i_{d}[t-n_{d}+1]]^{T}.$
Hence, at time $t$ the state variable $\bm{v}_{d}$ is a vector with entries equal to the last $n$ inputs to the system parametrized by the parameters $a_{i}$, where $n$ is the state dimension.
Note that although the $a_{i}$ parameters are free in general, in this paper we set $a_{i}=1$ for all $i$. Under this choice, the delay component $\bm{v}_{d}$ reduces to a simple temporal memory of the input, i.e.,
$\bm{v}_{d}[t]=[i_{d}[t],i_{d}[t-1],i_{d}[t-2],...,i_{d}[t-n_{d}+1]]^{T}.$ (5)
IV-B Proposed Model
Let $\bm{v}_{s}$ be the state variable of the general spiking neuron in (1) with parameter matrices $\bm{A}_{s},\bm{B}_{s},\bm{C}_{s}$ and state dimension $n_{s}$. Let $\bm{v}_{d}$ be the state variable of the time-shift SSM in Section IV-A2 with parameter matrices $\bm{A}_{d},\bm{B}_{d},\bm{C}_{d}$ and state dimension $n_{d}$. Our proposed spiking neuron model with state-delays is defined as:
$\bm{v}_{d}[t+1]$ $=\bm{A}_{d}\bm{v}_{d}[t]+\bm{B}_{d}i_{d}[t]$ (6a)
$\bm{v}_{s}[t+1]$ $=\bm{A}_{s}\bm{v}_{s}[t]+\bm{B}_{s}i_{s}[t]+\bm{A}_{sd}\bm{v}_{d}[t]-\bm{R}s[t]$ (6b)
$s[t]$ $=f_{\bf{\Theta}}(\bm{C}_{s}\bm{v}_{s}[t]+\bm{C}_{d}\bm{v}_{d}[t])$
$=\begin{cases}1\text{ if }(\bm{C}_{s}\bm{v}_{s}[t]+\bm{C}_{d}\bm{v}_{d}[t])\in\bf{\Theta}\\ 0\text{ otherwise}\end{cases}$ (6c)
Here, $\bm{A}_{sd}\in\mathbb{R}^{n_{s}\times n_{d}}$ gives the parameters that delayed inputs in the state $\bm{v}_{d}$ are multiplied with before they are added to the neuron state $\bm{v}_{s}$.
The above equations present a spiking neuron with additional access to temporal memory through the state variable $\bm{v}_{d}[t]$. The state dimension $n_{d}$ defines the maximum order of the delays provided to the spiking neuron i.e. the number of directly accessible past quantities. The input $i_{s}$ to $\bm{v}_{s}$, as defined in Section IV-A1, represents the input spikes to the neuron. The input $i_{d}$ to $\bm{v}_{d}$ represents the input stored in the temporal memory. See Table I for an overview of the notation. Examples of the proposed models with LIF and adLIF neuron models are provided in Sections IV-B1 and IV-B2, respectively. A high-level visualization of the proposed model is presented in Figure 2.
TABLE I: Table of notation.
Notation Meaning
$\bm{v}_{s}[t]$ State of the neuron’s intrinsic components
$\bm{v}_{d}[t]$ State of the neuron’s delay component
$n_{s}$ Dimension of $\bm{v}_{s}$
$n_{d}$ Dimension of $\bm{v}_{d}$ (delay order)
$\bm{A}_{s},\bm{B}_{s},\bm{C}_{s}$ Linear model parameters of $\bm{v}_{s}$
$\bm{A}_{d},\bm{B}_{d},\bm{C}_{d}$ Linear model parameters of $\bm{v}_{d}$
$\bm{R}$ Reset parameters
$\bm{A}_{sd}$ Delay parameters
$i_{s}[t]$ Input to $\bm{v}_{s}$
$i_{d}[t]$ Input to $\bm{v}_{d}$
$s[t]$ Spike output of the neuron
$l$ Number of hidden layers
$h$ Number of neurons in a hidden layer
Figure 2: High-level illustration of the proposed neuron model. The figure shows how the delay components are incorporated into the traditional SNN neuron model (LIF neuron shown).
IV-B1 LIF and RLIF example
The discrete-time Leaky Integrate-and-Fire (LIF) neuron, as defined in [bittar2022surrogate], with the additional proposed delays is given by:
$\bm{v}_{d}[t+1]$ $=\bm{A}_{d}\bm{v}_{d}[t]+\bm{B}_{d}i_{d}[t]$ (7a)
$\!u[t+1]$ $=\alpha u[t]+(1-\alpha)i_{s}[t]-\alpha\theta s[t]+\bm{A}_{sd}\bm{v}_{d}[t]$ (7b)
$s[t]$ $=f_{\theta}(u[t])=\begin{cases}1\text{ if }u[t]\geq\theta\\ 0\text{ otherwise.}\end{cases}$ (7c)
where $i_{s}[t]=\bm{W}\times\bm{s}^{l-1}[t]+\bm{V}\times\bm{s}^{l}[t]$ for RLIF and $i_{s}[t]=\bm{W}\times\bm{s}^{l-1}[t]$ for LIF. For both RLIF and LIF, we have $i_{d}[t]=\bm{W}\times\bm{s}^{l-1}[t]$.
IV-B2 adLIF and RadLIF example
The Adaptive Leaky Integrate-and-Fire (adLIF) neuron is an extension of the LIF neuron with an additional recovery state variable, $w[t]$ [bittar2022surrogate]. Discrete-time adLIF neuron with the additional proposed delays is:
$\bm{v}_{d}[t+1]=$ $\bm{A}_{d}\bm{v}_{d}[t]+\bm{B}_{d}i_{d}[t]$ (8a)
$\!u[t+1]=$ $\alpha u[t]+(1-\alpha)[i_{s}[t]-w[t]]-\alpha\theta s[t]$
$+\bm{A}_{sd}\bm{v}_{d}[t+1]$ (8b)
$\!w[t+1]=$ $au[t]+\beta w[t]+bs[t]$ (8c)
$s[t]=$ $f_{\theta}(u[t])=\begin{cases}1\text{ if }u[t]\geq\theta\\ 0\text{ otherwise.}\end{cases}$ (8d)
The inputs $i_{s}[t]$ and $i_{d}[t]$ are as defined for RLIF in Section IV-B1 with a correspondence between LIF and adLIF, and RLIF and RadLIF.
IV-C Delay parameters
We now provide an overview of all parameters related to the additional delay variable $\bm{v}_{d}$ in (6). We have $\bm{A}_{d}$ given by (3) under $a_{i}=1$ and $\bm{B}_{d}$ given by (4), hence they are fixed and non-trainable. Under LIF/RLIF neuron models of (7) and adLIF/RadLIF neuron models of (8), we do not use $\bm{C}_{d}$. Hence, the only additional (possibly trainable) parameter is $\bm{A}_{sd}$. In the rest of this section, we focus on $\bm{A}_{sd}$.
The matrix of parameters $\bm{A}_{sd}$ plays a central role in our development, as it determines the contribution of the delay variable $\bm{v}_{d}$ when added to the intrinsic neuron state $\bm{v}_{s}$, see (6). We consider four different settings for the elements of $\bm{A}_{sd}$:
  1. 1.
    $\bm{A}_{sd}=$ Ones, where $\bm{A}_{sd}]_{ij}=1$
  2. 2.
    $\bm{A}_{sd}=$ Linear decay, where $\bm{A}_{sd}]_{ij}=(n_{d}-j)\frac{1}{n_{d}}$
  3. 3.
    $\bm{A}_{sd}=$ Exponential decay, where $\bm{A}_{sd}]_{ij}=e^{-0.5j}$
  4. 4.
    $\bm{A}_{sd}\sim$ Uniform distribution, where $\bm{A}_{sd}]_{ij}\sim U(0,1]$
The equal sign == is used in the cases where the elements of $\bm{A}_{sd}$ are deterministic, while the sign $\sim$ is used in the case where the elements of $\bm{A}_{sd}$ are random. The matrix of delay parameters $\bm{A}_{sd}$ can be trainable or fixed. When it is fixed, its elements do not change during the training of the SNN and remain as initially set. When it is trainable, the initial set values of $\bm{A}_{sd}$ represent their initialization, and they change during the training phase without the constraint of preserving the distribution their initialization had, i.e., for instance if $\bm{A}_{sd}]_{ij}$ is initialized as ’Ones’, the parameters do not stay all ones or all equal. In Section VI we distinguish these cases by stating whether $\bm{A}_{sd}$ is non-trainable or trainable.
We now illustrate how the matrix $\bm{A}_{sd}$ determines the contribution of the delayed component $\bm{v}_{d}[t]$ in (5). The contributing delay component $\bm{A}_{sd}\bm{v}_{d}[t]$ is simply a weighted average of the past $n_{d}$ inputs $i_{d}[t]$. In particular, we have the following cases:
  1. 1.
    For $\bm{A}_{sd}=$ Ones:
    $\bm{A}_{sd}\bm{v}_{d}[t]=i_{d}[t]+i_{d}[t-1]+i_{d}[t-2]+$...
  2. 2.
    For $\bm{A}_{sd}=$ Linear Decay: For instance, with $n_{d}=5$:
    $\bm{A}_{sd}\bm{v}_{d}[t]=i_{d}[t]+0.8i_{d}[t-1]+0.6i_{d}[t-2]+$...
  3. 3.
    For $\bm{A}_{sd}=$ Exponential Decay: For instance, with $n_{d}=5$:
    $\bm{A}_{sd}\bm{v}_{d}[t]=i_{d}[t]+0.60i_{d}[t-1]+0.34i_{d}[t-2]+$...
  4. 4.
    For $\bm{A}_{sd}\sim$ Uniform. Here, the elements are random. A particular realization could be:
    $\bm{A}_{sd}\bm{v}_{d}[t]=0.1i_{d}[t]+0.9i_{d}[t-1]+0.2i_{d}[t-2]+$...
IV-D Computational complexity
The proposed neuron model with delays, given in (6), extends the baseline formulation in (1) by introducing an additional state variable $\bm{v}_{d}$ and its associated parameters $\bm{A}_{d}$, $\bm{B}_{d}$, $\bm{A}_{sd}$, and $\bm{C}_{d}$. The dimensionality of these matrices is determined by the delay order $n_{d}$. In this section, we analyze how this additional state and its parameters influence the computational complexity of the spiking neuron for the case of LIF and adLIF neuron as defined in Section IV-B1 and IV-B2.
IV-D1 Number of trainable parameters
The baseline definition of the LIF neuron model has $1$ trainable parameter, i.e., $\alpha$, while the adLIF neuron model has $4$ trainable parameters, i.e., $\alpha,\beta,a,b$. Let $l$ be the number of hidden layers, each with $h$ number of neurons. Let $c_{in}$ and $c_{out}$ be the number of input and output channels to the network. Then, the number of trainable parameters introduced by individual components of the networks is given by:
  • feedforward synaptic connections: $c_{in}h+h^{2}(l-1)+hc_{out}$
  • recurrent synaptic connections: $l(h^{2}-h)$
  • neuron parameters: $lh$ for LIF and =4lh=4lh for adLIF
  • normalization layers: $2(hl+c_{out})$
  • delays: $n_{d}hl$ if $\bm{A}_{sd}$ trainable, 0 otherwise.
From the above calculations, we observe that the number of additional parameters introduced by the delay mechanism is  $n_{d}hl$, and only if the matrix $\bm{A}_{sd}$ is trainable. Even in this case, the number of delay-related parameters remains relatively small compared to the number of synaptic connections, which scale as $h^{2}$. Relatively small values of $n_{d}$ get competitive performance, see Section. VI. Hence, since $h$ is much larger than both  $n_{d}$ or $l$ in both LIF and adLIF networks, the added computational and parametric overhead of incorporating delays is negligible.
In the remainder of this section, we compare the number of trainable parameters introduced by our proposed delay modeling with those reported in related work.
In [deckers2024co, Table 1], the inclusion of delays increased the total number of parameters from 38.738.7k to 75.875.8k i.e., by 37.137.1k parameters. For the same network architecture we increase the total number of parameters by $n_{d}hl=n_{d}\times 128\times 2$ parameters. For $n_{d}=10$, this corresponds to approximately 2.62.6k parameters, which is an order of magnitude smaller than in [deckers2024co].
[sun2025exploitingheterogeneousdelaysefficient, Section 2.1] reports that each neuron maintains a single learned delay shared across all outgoing connections, resulting in $\mathcal{O}(h)$ parameter scaling. In our formulation, up to $n_{d}$ delay-related parameters can be trained per neuron, where the overall scaling remains $\mathcal{O}(h)$. As noted in [sun2025exploitingheterogeneousdelaysefficient], this is considerably more efficient than the $\mathcal{O}(h^{2})$ scaling of synaptic weights.
IV-D2 Memory
The state variable $\bm{v}_{d}\in\mathbb{R}^{n_{d}\times 1}$ can be seen as a memory buffer that stores $n_{d}$ values per neuron.
The intrinsic neuron state $\bm{v}_{s}\in\mathbb{R}^{n_{s}\times 1}$ already contains $n_{s}$ internal variables that evolve over time. For example, the standard LIF neuron has a single state variable, the membrane potential, resulting in $n_{s}=1$. The adLIF, compared to the LIF, includes an additional threshold–adaptation variable, resulting in $n_{s}=2$. When the delay variable $\bm{v}_{d}$ is appended to the intrinsic state $\bm{v}_{s}$ in (6), the number of state variables for each neuron increases from $n_{s}$ to $n_{s}+n_{d}$.
At the network level, this increase applies to all hidden layers $l$ with $h$ neurons, increasing the total number of stored state variables from $n_{s}hl$ to $(n_{s}+n_{d})hl$. However, as noted in Section IV-D1, the values $n_{s},n_{d}$ and $l$ are all much smaller than $h$. Consequently, in both cases, with or without delays, the memory footprint scales predominantly with $h$.
V Experimental Set-up
V-1 Network Architecture and Training
In all experiments, a network with two hidden layers is used with batch normalization layers [ioffe2015batchnormalizationacceleratingdeep] and an accumulative output layer with cross entropy loss. Each model is trained using BPTT with surrogate gradient [neftci2019surrogate], AdamW optimizer [loshchilov2019decoupledweightdecayregularization] and cosine learning rate scheduler. We present further details for training and network parameters in Section IX-A.
V-2 Dataset
The Spiking Heidelberg Digits (SHD) dataset was introduced to standardize the evaluation of neuromorphic learning models [shd_ssc_dataset]. It consists of spike sequences derived from microphone recordings using the Lauscher artificial cochlea model. The dataset contains $20$ classes of spoken digits 0 to $9$ in both German and English. It has $8156$ training and $2264$ test samples. Each sample is represented by $700$ input channels. This dataset is primarily used for classification tasks, where the goal is to identify the spoken digit from the input spike sequence. Further details on dataset preprocessing and usage are provided in Section IX-B.
VI Results
In this section, we discuss the numerical experiments presented in Tables II, III, IV, V. Specifically in Section VI-A we comment on the performance of our proposed method in comparison to the works in the literature. In Sections VI-B, VI-C, VI-D, and VI-E we comment on the varying order of delays, the delay parameters, the corresponding network training time, and the varying number of neurons used, respectively. In the final Section VI-F we discuss the effect of trainable and non-trainable delay parameters.
TABLE II: Accuracy and number of parameters reported for delay-based SNN papers on the SHD dataset. M denotes millions of parameters.
Paper Accuracy # params.
Learning DCLS [hammouamri2023learning] 95.07 $\pm$ 0.24 % 0.2M
Co-learning [deckers2024co] 95.02 % 0.076M
Learnable axonal delay [10.3389/fnins.2023.1275944] 93.55 % 0.21M
Event-based Delay [meszaros2025efficient] 93.24 $\pm$ 1.00 % 0.5M (a)
Temporal Hierarchy [moro2024roletemporalhierarchyspiking] 92.1 % (b) -
Heterogeneous Delays [sun2025exploitingheterogeneousdelaysefficient] 90.98 % 0.1M (c)
Our (adLIF, $n_{d}=5$, non-train. $\bm{A}_{sd}$) 94.1 $\pm$ 0.5 % 0.04M
  • (a)
    Read from [meszaros2025efficient, Fig. 5]
  • (b)
    Read from [moro2024roletemporalhierarchyspiking, Fig. 5]
  • (c)
    Read from [Fig.1c] [sun2025exploitingheterogeneousdelaysefficient]
TABLE III: Test accuracy on the SHD dataset across combinations of neuron model, delay parameters $\bm{A}_{sd}$, and delay order $n_{d}$. All networks consist of two hidden layers with $h=128$ neurons each. Bold values indicate the best performance for each neuron model (per row).
(a) $\bm{A}_{sd}$ is non-trainable.
Neuron Model No delay $\bm{A}_{sd}=$ Ones $\bm{A}_{sd}=$ Lin. Decay $\bm{A}_{sd}=$ Exp. Decay $\bm{A}_{sd}\sim$ Uniform
$n_{d}=5$ $n_{d}=10$ $n_{d}=5$ $n_{d}=10$ $n_{d}=5$ $n_{d}=10$ $n_{d}=5$ $n_{d}=10$
LIF 85.9 $\pm$ 1.2 % 86.6 $\pm$ 1.2 % 83.8 $\pm$ 2.0 % 86.6 $\pm$ 0.8 % 85.9 $\pm$ 0.8 % 86.7 $\pm$ 0.7 % 86.0 $\pm$ 0.9 % 87.2 $\pm$ 1.6 % 86.7 $\pm$ 0.8 %
RLIF 90.6 $\pm$ 0.5 % 87.7 $\pm$ 1.1 % 87.1 $\pm$ 1.1 % 88.8 $\pm$ 1.0 % 87.8 $\pm$ 0.6 % 87.9 $\pm$ 1.7 % 89.0 $\pm$ 0.7 % 89.6 $\pm$ 1.0 % 88.0 $\pm$ 0.9 %
adLIF 92.0 $\pm$ 0.4 % 92.1 $\pm$ 0.2 % 91.6 $\pm$ 0.4 % 93.3 $\pm$ 0.9 % 91.8 $\pm$ 0.8 % 93.3 $\pm$ 0.4 % 93.0 $\pm$ 0.9 % 94.1 $\pm$ 0.5 % 92.1 $\pm$ 1.3 %
RadLIF 93.4 $\pm$ 1.1 % 91.6 $\pm$ 2.6 % 91.2 $\pm$ 2.2 % 92.8 $\pm$ 0.7 % 92.4 $\pm$ 0.9 % 93.8 $\pm$ 0.6 % 93.5 $\pm$ 0.4 % 93.1 $\pm$ 1.2 % 92.1 $\pm$ 1.2 %
(b) $\bm{A}_{sd}$ is trainable.
Neuron Model $\bm{A}_{sd}=$ Ones init. $\bm{A}_{sd}=$ Lin. Decay init. $\bm{A}_{sd}=$ Exp. Decay init. $\bm{A}_{sd}\sim$ Uniform init.
$n_{d}=5$ $n_{d}=10$ $n_{d}=5$ $n_{d}=10$ $n_{d}=5$ $n_{d}=10$ $n_{d}=5$ $n_{d}=10$
LIF 87.7 $\pm$ 0.9 % 88.2 $\pm$ 1.0 % 88.4 $\pm$ 1.2 % 89.1 $\pm$ 0.3 % 89.1 $\pm$ 0.8 % 89.6 $\pm$ 1.1 % 88.4 $\pm$ 1.5 % 89.5 $\pm$ 0.5 %
RLIF 88.5 $\pm$ 1.7 % 87.9 $\pm$ 1.3 % 89.0 $\pm$ 0.5 % 89.5 $\pm$ 1.2 % 89.1 $\pm$ 0.6 % 90.3 $\pm$ 0.9 % 88.9 $\pm$ 0.5 % 89.7 $\pm$ 0.8 %
adLIF 92.7 $\pm$ 0.7 % 91.9 $\pm$ 0.5 % 93.1 $\pm$ 0.4 % 92.7 $\pm$ 0.5 % 93.6 $\pm$ 0.2 % 93.6 $\pm$ 1.1 % 93.3 $\pm$ 0.6 % 92.6 $\pm$ 0.5 %
RadLIF 93.4 $\pm$ 0.8 % 92.4 $\pm$ 1.3 % 94.0 $\pm$ 0.2 % 93.3 $\pm$ 0.6 % 93.4 $\pm$ 0.5 % 93.7 $\pm$ 0.6 % 93.7 $\pm$ 0.6 % 92.3 $\pm$ 1.0 %
VI-A Comparison to literature
In this section, we compare our best performance presented in this paper with the performances of SHD dataset with other SNN works which include delays, see Table II. We observe that our model is comparable both in terms of performance and in terms of number of parameters used.
Note that in general, the highest average accuracy reported over multiple initialization instances for the SHD dataset using SNNs is $95.07\%$ [hammouamri2023learning], which is included in Table II. The benchmark performance using a non-spiking encoding of the same data and a recurrent neural network is $99.69\%$ [boeshertz2024accuratemappingrnnsneuromorphic].
VI-B Order of delays
In this section, we discuss the varying order of delays i.e., we compare $n_{d}=0$, $5$ and $10$, see Table III.
We observe that increasing the delay order from $n_{d}=0$ (i.e. no delay, hence the same as the traditional neuron types) to $n_{d}=5$ or $n_{d}=10$ in the LIF, adLIF, and RadLIF neurons either maintains (i.e. the performance with and without delay are within standard deviations of each other) or improves performance. For instance, in Table III(a) for the adLIF neuron model, the baseline accuracy with $n_{d}=0$ is $92.0\pm 0.4\%$, while for $\bm{A}_{sd}=\text{Uniform}$ and $n_{d}=5$ the accuracy rises to $94.1\pm 0.5\%$, indicating an improvement beyond the standard deviation range. In contrast, the RLIF neuron either maintains its performance or shows only a slight decrease when delays are introduced. For example, in Table III(a) for the RLIF neuron model, the baseline accuracy with $n_{d}=0$ is $90.6\pm 0.5\%$, while for $\bm{A}_{sd}=\text{Ones}$ and $n_{d}=5$ the accuracy drops to $87.7\pm 1.1\%$, indicating a performance drop beyond the standard deviation range.
When comparing $n_{d}=5$ and $n_{d}=10$, the results remain within the standard deviation of each other, suggesting that increasing the delay order to $n_{d}=10$ does not yield additional benefits beyond what $n_{d}=5$ already captures.
VI-C Delay parameters
In this section we investigate the impact of using various delay parameter settings for $\bm{A}_{sd}$. In particular, we explore $\bm{A}_{sd}=$ Ones, Linear Decay, Exponential Decay, and $\bm{A}_{sd}\sim$ Uniform, see Section IV-C for more details of the setting. Table III provides the associated results.
We do not observe a consistent trend that holds in all values of $n_{d}$ and in all neuron types. Most of the accuracy values are within the standard deviation of each other. The remaining differences appear to be model-specific. For instance, considering the average performance of the adLIF neuron, the best configuration for a non-trainable $\bm{A}_{sd}$ is the Uniform setting, whereas for a trainable $\bm{A}_{sd}$, the Exponential Decay initialization yields the highest accuracy.
VI-D Training time
In this section, we discuss the additional training time introduced by using delays, see Table IV. We first focus on the increase in training time when using $n_{d}=5$ compared to the baseline with no delay. The smallest ratio increase, $\frac{13.1}{8.6}=1.5$, occurs for the adLIF model with a non-trainable $\bm{A}_{sd}$, while the largest, $\frac{12.5}{6.5}=1.8$, is observed for the LIF model with a trainable $\bm{A}_{sd}$. Increasing $n_{d}$ from $5$ to $10$ results in only a negligible change in training time.
TABLE IV: Training time in minutes as average over the various delay parameters $\bm{A}_{sd}$ settings for fixed $n_{d}$ from Table III.
(a) $\bm{A}_{sd}$ non-trainable
Neuron Model No delay $n_{d}=5$ $n_{d}=10$
LIF 6.5 $\pm$ 0.1 min. 11.7 $\pm$ 0.1 min. 12.0 $\pm$ 0.2 min.
RLIF 7.9 $\pm$ 0.1 min. 13.1 $\pm$ 0.2 min. 12.9 $\pm$ 0.1 min.
adLIF 8.6 $\pm$ 0.1 min. 13.1 $\pm$ 0.3 min. 13.4 $\pm$ 0.6 min.
RadLIF 9.6 $\pm$ 0.1 min. 15.4 $\pm$ 0.2 min. 15.1 $\pm$ 0.1 min.
(b) $\bm{A}_{sd}$ trainable
Neuron Model $n_{d}=5$ $n_{d}=10$
LIF 12.5 $\pm$ 0.2 min. 12.7 $\pm$ 0.3 min.
RLIF 13.9 $\pm$ 0.2 min. 13.7 $\pm$ 0.1 min.
adLIF 14.1 $\pm$ 0.2 min. 13.9 $\pm$ 0.1 min.
RadLIF 15.9 $\pm$ 0.4 min. 16.0 $\pm$ 0.3 min.
VI-E Number of neurons vs order of delays
In this section, we discuss the combined effect of the number of neurons in the hidden layers $h$ and the order of delays $n_{d}$, see Table V.
The impact of increasing $n_{d}$ on the performance depends on the number of neurons $h$. For $h\geq 32$, using delays ($n_{d}=10,50$ or $100$) maintains performance at a level similar to no-delay or slightly reduces it. In contrast, for smaller networks with $h\leq 16$, using delays yields a clear improvement over the baseline no-delay performance.
It is worth noting that, in our implementation, the input sequences of the SHD dataset have a length of $100$. Consequently, setting $n_{d}=100$ allows the network to access the entire temporal history of each sample at every time step.
TABLE V: Test accuracy on the SHD dataset across combinations of number of neurons $h$ and delay order $n_{d}$. All networks consist of two hidden layers with $h$ neurons each, adLIF neuron model and $\bm{A}_{sd}\sim$ Uniform used as a fixed value in the $\bm{A}_{sd}$ non-trainable case and as initialization in the $\bm{A}_{sd}$ trainable case. Bold values indicate the best performance for each $h$ value (per row).
(a) $\bm{A}_{sd}$ non-trainable.
$h$ No delay $n_{d}=10$ $n_{d}=50$ $n_{d}=100$
512 92.5 $\pm$ 0.7 % 92.7 $\pm$ 0.9 % 90.8 $\pm$ 1.5 % 90.4 $\pm$ 1.0 %
256 93.1 $\pm$ 1.1 % 92.1 $\pm$ 0.5 % 90.6 $\pm$ 1.3 % 89.9 $\pm$ 0.6 %
128 92.0 $\pm$ 0.4 % 92.1 $\pm$ 1.3 % 89.8 $\pm$ 1.2 % 89.4 $\pm$ 0.6 %
64 91.3 $\pm$ 1.1 % 91.8 $\pm$ 1.3 % 89.6 $\pm$ 1.1 % 89.4 $\pm$ 1.8 %
32 88.7 $\pm$ 2.2 % 90.2 $\pm$ 1.3 % 86.8 $\pm$ 1.4 % 88.2 $\pm$ 1.2 %
16 78.1 $\pm$ 1.0 % 86.9 $\pm$ 1.2 % 86.6 $\pm$ 0.7 % 87.1 $\pm$ 1.2 %
8 60.1 $\pm$ 7.4 % 75.3 $\pm$ 1.6 % 77.3 $\pm$ 2.2 % 80.3 $\pm$ 1.3 %
(b) $\bm{A}_{sd}$ trainable.
$h$ $n_{d}=10$ $n_{d}=50$ $n_{d}=100$
512 92.7 $\pm$ 0.9 % 90.8 $\pm$ 1.5 % 90.4 $\pm$ 1.0 %
256 92.1 $\pm$ 0.5 % 90.6 $\pm$ 1.3 % 89.9 $\pm$ 0.6 %
128 92.1 $\pm$ 1.3 % 89.8 $\pm$ 1.2 % 89.4 $\pm$ 0.6 %
64 91.8 $\pm$ 1.3 % 89.6 $\pm$ 1.1 % 89.4 $\pm$ 1.8 %
32 90.2 $\pm$ 1.3 % 86.8 $\pm$ 1.4 % 88.2 $\pm$ 1.2 %
16 86.2 $\pm$ 1.0 % 84.1 $\pm$ 1.2 % 83.0 $\pm$ 1.0 %
8 72.6 $\pm$ 2.5 % 70.1 $\pm$ 2.6 % 73.8 $\pm$ 1.9 %
VI-F Trainable vs non-trainable delay parameters
In this section, we compare the results obtained with trainable and non-trainable $\bm{A}_{sd}$ matrices.
Comparing Tables III(a) and III(b), we observe that the results are generally within the standard deviation of each other, except for the LIF case, where the trainable $\bm{A}_{sd}$ provides a clear performance benefit. Similarly, Tables V(a) and V(b) show no meaningful performance difference between the two configurations.
Tables IV(a) and IV(b) indicate that the additional training time required when $\bm{A}_{sd}$ is trainable is negligible <1<1 minute, which further supports the discussion in Section IV-D1 regarding the small computational complexity overhead when using our proposed delays.
VII Discussions
The results suggest that the LIF and adLIF neurons may not be fully capable of exploiting the introduced delay mechanism. Their limited state dimensionality and small number of free parameters likely restrict the model’s ability to fully utilize temporal depth introduced by the delay mechanism. More complex neuron models with richer internal dynamics might benefit more from the proposed delay formulation.
We did not perform any dedicated hyperparameter optimization (HPO) in this study. Conducting a thorough HPO for each neuron type and delay configuration could further improve performance and potentially alter observations.
From a neuromorphic computing perspective, an important open question is which of the proposed delay neuron models can be implemented efficiently on hardware. Implementing delays on hardware has been explored [10181778, DelGrad, meijer2025efficientsynapticdelayimplementation, DenRAM], programmable delays can be implemented on various hardware including Intel Loihi [8259423], IBM TrueNorth [7229264], SpiNNaker [6750072], and SENeCA [9870025]. Whether similar implementations can be performed for the proposed models is an open question.
VIII Conclusion
In this work, we proposed a general framework for modeling delays within spiking neuron dynamics by introducing additional state variables that encodes delayed information. This formulation enables neurons to access a finite temporal history. The framework is general and can be easily integrated into various spiking neuron models.
We analyzed different aspects of this formulation, including the order of delays, delay parameters and their optimization (trainable or fixed), and the trade-off between delay order and network size.
Experimental results show that the proposed delay mechanism achieves performance comparable to that of existing delay-based approaches while remaining computationally efficient. The proposed delay mechanism introduces either no additional trainable parameters or only a small number relative to the total network size. Furthermore, we observe that incorporating delays significantly improves performance in smaller networks, where the delays can compensate for the restricted inherent capacity of the model due to the limited number of neurons.
IX Appendix
IX-A Network and Training further details
Our code is based on the Sparch implementation [bittar2022surrogate]. During the training process,neuron parameters are clipped to remain within these boundaries as in [deckers2024co] i.e. $\alpha\in[0.36,0.96],\beta\in[0.96,0.99],a\in[0,1]$ and $b\in[0,2]$.
We use hyperparameters based on previous works [hammouamri2023learning, deckers2024co, schöne2024scalableeventbyeventprocessingneuromorphic], in particular base LR $10^{-2}$, WD $10^{-5}$, dropout rate =0.4=0.4, batch size =128=128, and number of epochs =50=50. No hyperparameter optimization has been performed.
IX-B Dataset further details
For the SHD dataset, similar to other works [hammouamri2023learning] [deckers2024co] we used spatio-temporal bins to reduce the input dimensions. The input channels were reduced from $700$ to $140$ by binning every $5$ consecutive channels.
The dataset was put in frames using the To-Frame [tonic] function with time-window=10000=10000, i.e. accumulation of events with time-windows of 1010ms. The framed data set was used to load batches during both training and testing. During training, we use the data augmentation techniques TimeNeurons_mask_aug and CutMix following previous work [deckers2024co] [hammouamri2023learning].

### v2

@@
 Delays in Spiking Neural Networks: A State Space Model Approach
-Delays in Spiking Neural Networks:
-A State Space Model Approach
-Sanja Karilanova, Subhrakanti Dey, Ayça Özçelikkale
-Department of Electrical Engineering, Uppsala University, Sweden
-{Sanja.Karilanova, Subhrakanti.Dey, Ayca.Ozcelikkale}@angstrom.uu.se
-S.
-Karilanova and A.
-Özçelikkale acknowledge the support of Center for Interdisciplinary Mathematics (CIM), and AI4Research, Uppsala University.The computations were enabled by resources provided by the National Academic Infrastructure for Supercomputing in Sweden (NAISS), partially funded by the Swedish Research Council through grant agreement no.
-2022-06725.
-I Abstract
-Spiking neural networks (SNNs) are biologically inspired, event-driven models that are suitable for processing temporal data and offer energy-efficient computation when implemented on neuromorphic hardware.
+Abstract
+Spiking neural networks (SNNs) are biologically inspired, event-driven models suited for temporal data processing and energy-efficient neuromorphic computing.
 In SNNs, richer neuronal dynamic allows capturing more complex temporal dependencies, with delays playing a crucial role by allowing past inputs to directly influence present spiking behavior.
@@
 The proposed mechanism enables each neuron to access a finite temporal input history.
-The framework is agnostic to neuron models and hence can be seamlessly integrated into standard spiking neuron models such as LIF and adLIF.
+The framework is agnostic to neuron models and hence can be seamlessly integrated into standard spiking neuron models such as Leaky Integrate-and-Fire (LIF) and Adaptive LIF (adLIF).
 We analyze how the duration of the delays and the learnable parameters associated with them affect the performance.
 We investigate the trade-offs in the network architecture due to additional state variables introduced by the delay mechanism.
-Experiments on the Spiking Heidelberg Digits (SHD) dataset show that the proposed mechanism matches the performance of existing delay-based SNNs while remaining computationally efficient.
-Moreover, the results illustrate that the incorporation of delays may substantially improve performance in smaller networks.
+Experiments on the Spiking Heidelberg Digits (SHD) dataset show that the proposed mechanism matches existing delay-based SNNs in performance while remaining computationally efficient, with particular gains in smaller networks.
 Code available at https://github.com/sannkka/Modeling-State-Delays-in-SNNs.
-II Introduction
-Spiking Neural Networks (SNNs) [MAASS19971659] have emerged as a more biologically plausible class of neural models compared to conventional artificial neural networks (ANNs), offering promising advantages in temporal processing and energy efficiency, particularly when deployed on neuromorphic hardware [davies_advancing_2021].
-SNNs encode information in the timing of discrete spikes, allowing them to capture temporal structure and making them well suited for modeling time-dependent signals such as audio [bittar2022surrogate], sensory [10.3389/fnins.2022.951164], or visual patterns [Gallego2022].
+1 Introduction
+Spiking Neural Networks (SNNs) have emerged as a more biologically plausible class of neural models compared to conventional artificial neural networks (ANNs), offering promising advantages in temporal processing and energy efficiency, particularly when deployed on neuromorphic hardware [6].
+SNNs encode information in the timing of discrete spikes, allowing them to capture temporal structure and making them well suited for modeling time-dependent signals such as audio [1], or visual patterns [13].
 Despite these advantages, achieving performance comparable to that of ANNs remains an open challenge, in part due to the limited temporal expressivity of current SNN models.
 In biological neural systems, temporal delays play a critical computational role.
-Delays that enable neurons to integrate information from multiple past moments support tasks such as hearing [carr1993processing] and brain activity [carr1990circuit].
+Delays that enable neurons to integrate information from multiple past moments support tasks such as hearing [3] and brain activity [4].
 Inspired by these findings, numerous works have explored incorporating delays into computational SNNs models.
-Theoretical analyses have shown that an SNN with $k$ adjustable delays can compute a much richer class of functions than a threshold circuit with $k$ adjustable weights [MAASS199926].
-Empirical studies have also demonstrated improved performance when optimizing spike transmission delays along with synaptic weights or membrane time constants [deckers2024co, sun2025exploitingheterogeneousdelaysefficient].
+Theoretical analyses have shown that an SNN with $k$ adjustable delays can compute a much richer class of functions than a threshold circuit with $k$ adjustable weights [24].
+Empirical studies have also demonstrated improved performance when optimizing spike transmission delays along with synaptic weights or membrane time constants [7, 32].
 These results suggest that delays can enrich the information processing capabilities of SNNs and their internal representation.
@@
 In this work, we propose a method based on State Space Models (SSMs) for equipping SNNs with delays.
-The method uses a time-shift based state transition, which enables each neuron to access both its current input and the most recent history of past input values.The formulation is general and can be incorporated into arbitrary spiking neuron models.
+The method uses a time-shift based state transition, which enables each neuron to access both its current input and the most recent history of past input values.
+The formulation is general and can be incorporated into arbitrary spiking neuron models, such as Leaky Integrate-and-Fire (LIF) and Adaptive LIF (adLIF) and their recurrent variant Recurrent LIF (RLIF) and Recurrent adLIF (RadLIF) [1].
 Conceptually, the framework bridges ideas from SSMs and SNNs, allowing SNNs to capture a fixed number of past inputs in a structured and interpretable manner.
 Our contributions can be summarized as follows:
-  • We propose a method for incorporating delays within the state dynamics through additional delay variables, enabling neurons to access a finite temporal history of inputs.
-  • The proposed formulation is general and easily integrable into existing spiking neuron models, including LIF, adLIF, and RadLIF.
-  • We conduct a systematic analysis of delay order, i.e.
-the number of allowed past inputs, and their associated learnable parameters, and study the trade-offs between the number of neurons in the network and the delay order.
-  • The proposed framework introduces either no additional trainable parameters (i.e when the delay parameters are not trained) or a number of trainable parameters comparable to the other SNN delay models.
+  • We propose incorporating delays within the neuron state dynamics via additional delay variables, enabling access to a finite temporal history of inputs.
+  • The formulation is general and easily integrable into existing spiking neuron models, including LIF, RLIF, adLIF, and RadLIF.
+  • We systematically analyze delay order (number of accessible past inputs) and associated learnable parameters, studying trade-offs between network size and delay order.
+  • The framework introduces either no additional trainable parameters (non-trainable case) or a parameter count comparable to other SNN delay models.
 The proposed models achieve performance comparable to existing delay-based SNN approaches on the Spiking Heidelberg Digits (SHD) dataset.
-Moreover, the observed trade-offs highlight promising directions for future exploration, particularly showing that incorporating delays can substantially improve accuracy in smaller SNNs where the neuron count limits temporal capacity.
-Figure 1: Illustration of different approaches of modeling delays. Our proposed delays are modeled within the state dynamics of the neuron (green). Figure is inspired by [DelGrad, Fig. 3(a)].
-III Related Work
-In the SNN literature, three main types of delays are commonly modeled: axonal delays, applied to the output of a neuron [shrestha2018slayer, 10.3389/fnins.2023.1275944, sun2025exploitingheterogeneousdelaysefficient]; dendritic delays, applied to its input [diaz2016efficient]; and synaptic delays, specific to each connection between neuron pairs [10.1145/3589737.3606009, deckers2024co, hammouamri2023learning, dominijanni2025extendingspiketimingdependentplasticity, meszaros2024learning].
+The observed trade-offs highlight promising directions for future exploration, particularly showing that incorporating delays can substantially improve accuracy in smaller SNNs where the neuron count limits temporal capacity.
+2 Related Work
+In the SNN literature, three main types of delays are commonly modeled: axonal delays, applied to the output of a neuron [31, 33, 32]; dendritic delays, applied to its input [8]; and synaptic delays, specific to each connection between neuron pairs [15, 7, 17, 9].
 In contrast to these formulations, we propose an SSM based approach where delays are modeled within the state dynamics of the neuron.
 See Figure 1 for a schematic illustration.
-The most common ways to model delays in SNNs is to learn a single delay value applied to each spike [shrestha2018slayer], or to employ temporal convolutions [hammouamri2023learning].
-Several works explore the role of delays in SNNs by jointly training weights and delays [deckers2024co], learning only delays [10.1145/3589737.3606009], or showing that heterogeneous delays can reduce weight-precision requirements [sun2025exploitingheterogeneousdelaysefficient].
-Others relate delays to temporal parameters such as time constants [10.1371/journal.pcbi.1012673] or study their role in the formation of temporal hierarchies between layers [moro2024roletemporalhierarchyspiking].
-In these works, various learning algorithms have been used, such as variations of backpropagation through time (BPTT) [hammouamri2023learning, shrestha2018slayer, queant2025delreclearningdelaysrecurrent, meszaros2024learning], biologically inspired spike-timing-dependent plasticity (STDP) [dominijanni2025extendingspiketimingdependentplasticity], and exact spike-time gradients [DelGrad, meszaros2025efficient].
-Our formulation represents delays as state variables evolving over time.
-This structure remains compatible with BPTT, enabling training with the popular surrogate-gradient methods [neftci2019surrogate] used to train SNNs.
+A typical approach for modeling delays in SNNs is through explicit temporal convolutions applied to spike trains [17, eqn.
+6–8], [33, eqn.
+4], [31, eqn.
+3], where a kernel with a relatively small number of nonzero weights is convolved with the spike train, and the relative positions of the non-zero weights encode the delay mechanism.
+A related approach, and the one closest to our proposed method, is to incorporate temporal buffer directly into the neuron state as an explicit finite memory of past inputs [34].
+Given that linear state-space models may be interpreted as temporal convolutions [12], this type of approaches can be seen as an extra incorporation of delay into temporal convolution, see also (3).
+An alternative line of work models delays as continuous-valued spike-time shifts [14, 25], which permits exact gradient methods but is less compatible with discrete surrogate-gradient training pipelines.
+Figure 1: Delay modeling approaches. Proposed method: delays within neuron state dynamics (green). Figure inspired by [14, Fig. 3(a)].
+The role of delays in SNNs has been studied from various perspectives including jointly training weights and delays [7], learning only delays [15], exploring heterogeneous delays and their relation to weight-precision requirements [32], and investigating their role in temporal hierarchies across layers [26].
+Learning algorithms for delays range from BPTT [17, 31] and STDP [9] to exact spike-time gradients [14, 25].
+Our formulation represents delays as state variables, making it compatible with BPTT and the widely used surrogate-gradient method for training SNNs [27].
 Delays have long been studied in fields such as signal processing and control theory.
-In discrete-time dynamical systems, they are typically modeled through delay-difference or state-space formulations, where past states are explicitly included to capture finite memory effects [RICHARD20031667, SenameOlivier2005Sots].
-Recently, modern deep SSMs, which stack multiple state-space layers combined with nonlinearities, have demonstrated strong performance in sequence modeling tasks [gu2022efficientlymodelinglongsequences, Gupta_DSS].
-Even more recently, parameterization has been introduced in deep SSMs, effectively capturing delayed state dependencies [fu2022hungry, zhang2023effectively].
-We build on this formulation to represent controllable delays within spiking neural network models.
-A common approach in modern SNNs embeds spiking dynamics within SSMs, where at each time step the output of the deep SSM layer is passed through a firing threshold function [stan2023learning, bal2024rethinkingspikingneuralnetworks, karilanova2025statespacemodelinspiredmultipleinput, karilanova_reset, du2024spiking].
-We continue in this direction by utilizing the SSM perspective on delay modeling to spiking neural architectures, enabling SNNs to incorporate structured temporal dependencies through the same principled state-space formulation.
-IV Methods
-IV-A Preliminaries
-IV-A1 General Spiking Neuron Model
-A single general feed-forward discrete-time $n$-state dimensional spiking neuron can be defined as [karilanova2025zeroshottemporalresolutiondomain, karilanova2025statespacemodelinspiredmultipleinput, karilanova_reset]:
+In discrete-time dynamical systems, they are typically modeled through delay-difference and state-space formulations [28, 30].
+Modern deep SSMs, which stack multiple state-space layers combined with nonlinearities, have demonstrated strong performance in sequence modeling tasks [16].
+Subsequent work introduced parameterization in deep SSMs, which effectively captures delayed state dependencies [11, 36].
+In parallel, hybrid approaches have emerged that bridge SNNs and SSMs, combining properties and capabilities from both models [35, 20, 19, 10, 21, 32].
+We build on this line of work and adopt the SSM perspective on delay modeling within spiking neural architectures.
+This allows SNNs to incorporate structured temporal dependencies through a principled state-space formulation.
+3 Methods
+3.1 Preliminaries
+3.1.1 General Spiking Neuron Model
+A single general feed-forward discrete-time $n$-state dimensional spiking neuron can be defined as [21]:
 $\!\bm{v}_{s}[t+1]$ $=\bm{A}_{s}\bm{v}_{s}[t]+\bm{B}_{s}i_{s}[t]-\bm{R}s[t]$ (1a)
@@
 The spiking behavior is described by $f_{\bf{\Theta}}(\cdot)$, where the neuron spikes when the state enters into the region described by $\bf{\Theta}$.
-The neuron model in (1) generalizes several commonly used formulations, including the Leaky Integrate-and-Fire (LIF) and adaptive LIF (adLIF) neurons, see [karilanova2025zeroshottemporalresolutiondomain, karilanova2025statespacemodelinspiredmultipleinput, karilanova_reset] for illustration of such representations.
-Spiking neural networks are built by connecting neurons such that the output spikes from the previous (pre-synaptic, l1l-1) layer serve as inputs to the next (post-synaptic, $l$) layer.
-Let $L_{\text{pre}}$ and $L_{\text{post}}$ be the number of neurons in the pre and post synaptic layer.
-Let $\bm{W}\in\mathbb{R}^{L_{\text{post}}\times L_{\text{pre}}}$ be the weight matrix connecting the two layers, and $\bm{V}\in\mathbb{R}^{L_{\text{post}}\times L_{\text{post}}}$ be the weight matrix associated with the recurrent connections between neurons in a layer.
-Then $\bm{i}[t]=\bm{W}\times\bm{s}^{l-1}[t]+\bm{V}\times\bm{s}^{l}[t]$, where $\bm{s}^{l-1}[t]$ and $\bm{s}^{l}[t]$ are a vector of all the spikes from the previous and current layer, respectively.
-IV-A2 State Space Model with Time Shift State Transition Matrix
-A discrete-time time-invariant linear SSM can be written as [ljung_SI]:
+The model in (1) generalizes commonly used formulations including including the Leaky Integrate-and-Fire (LIF) and adaptive LIF (adLIF) neurons and their recurrent variants Recurrent LIF (RLIF) and Recurrent adLIF (RadLIF) [21].
+In SNNs, output spikes from the pre-synaptic layer (l1l-1) serve as inputs to the post-synaptic layer ($l$).
+Let $\bm{W}\in\mathbb{R}^{L_{\text{post}}\times L_{\text{pre}}}$ and $\bm{V}\in\mathbb{R}^{L_{\text{post}}\times L_{\text{post}}}$ be the feedforward and recurrent weight matrices, with $L_{\text{pre}}$, $L_{\text{post}}$ the layer sizes.
+The input current is $\bm{i}[t]=\bm{W}\times\bm{s}^{l-1}[t]+\bm{V}\times\bm{s}^{l}[t]$, where $\bm{s}^{l-1}[t]$ and $\bm{s}^{l}[t]$ are spike vectors from the previous and current layer.
+3.1.2 Linear State Space Model
+A discrete-time time-invariant linear SSM can be written as [16]:
 $\bm{v}_{d}[t+1]=\bm{A}_{d}\bm{v}_{d}[t]+\bm{B}_{d}i_{d}[t]$ (2a)
@@
 The matrices $\bm{B}_{d}\in\mathbb{R}^{n_{d}\times 1},\bm{C}_{d}\in\mathbb{R}^{n_{out}\times n_{d}},\bm{D}_{d}\in\mathbb{R}^{n_{out}\times 1}$ describe the interaction of the SSM through its inputs and outputs.
-The lower shift matrix is a structured form of the state-transition matrix with nonzero elements only on the subdiagonal. Let $\bm{A}_{d}$ be the lower shift matrix:
-$\bm{A}_{d}$ $=\begin{bmatrix}0&0&\ldots&0&0\\ a_{1}&0&\ldots&0&0\\ 0&a_{2}&\ldots&0&0\\ \vdots&&\ddots&\vdots&\vdots\\ 0&0&\ldots&a_{n_{d}-1}&0\end{bmatrix}\text{i.e.,}$ (3a)
-$[\bm{A}_{d}]_{i,j}$ $=\begin{cases}a_{j}&\text{ for }i-1=j\\ 0&\text{ otherwise }\end{cases}$ (3b)
-where $\bm{A}_{d}]_{i,j}$ represents the $i$th row $j$th column element of the matrix $\bm{A}_{d}$. Let
-$\bm{B}_{d}=[1,0,\ldots,0]^{T}.$ (4)
-Then for the SSM state variable $\bm{v}_{d}$ defined in (2) and initial state $\bm{v}_{d}[0]=\bm{0}^{n_{d}\times 1}$ we have:
-$\bm{v}_{d}[1]$ $=[i_{d}[1],\,0,\,0,\,...,\,0]^{T}$
-$\bm{v}_{d}[2]$ $=[i_{d}[2],\,a_{1}i_{d}[1],\,0,\,...,\,0]^{T}$
-$\bm{v}_{d}[3]$ $=[i_{d}[3],\,a_{1}i_{d}[2],\,a_{2}a_{1}i_{d}[1],\,...,\,0]^{T}$
-$\,\,\,\vdots$
-$\bm{v}_{d}[t]$ $=[i_{d}[t],\,a_{1}i_{d}[t-1],\,a_{2}a_{1}i_{d}[t-2],\,$...,\,
-$\hskip 85.35826pt(\prod^{n-1}_{k=1}a_{k})i_{d}[t-n_{d}+1]]^{T}.$
-Hence, at time $t$ the state variable $\bm{v}_{d}$ is a vector with entries equal to the last $n$ inputs to the system parametrized by the parameters $a_{i}$, where $n$ is the state dimension.
-Note that although the $a_{i}$ parameters are free in general, in this paper we set $a_{i}=1$ for all $i$.
-Under this choice, the delay component $\bm{v}_{d}$ reduces to a simple temporal memory of the input, i.e.,
-$\bm{v}_{d}[t]=[i_{d}[t],i_{d}[t-1],i_{d}[t-2],...,i_{d}[t-n_{d}+1]]^{T}.$ (5)
-IV-B Proposed Model
+3.1.3 Temporal Convolution Representation of Linear SSM
+A key property of linear SSMs is that it can be expressed as a causal temporal convolution. Under the initial condition $v_{d}[t]=0$, for $t\leq 0$, unrolling the recurrence in (2a) yields
+$\bm{v}_{d}[t]=\sum_{k=0}^{t-1}\bm{A}_{d}^{k}\bm{B}_{d}\,i_{d}[t-k].$ (3)
+The state variable at time $t$ is therefore a convolution of the input with the kernel $d[t]=\bm{A}_{d}^{t}\bm{B}_{d}$ i.e. $\bm{v}_{d}[t]=d[t]*i_{d}[t]$, where * denotes convolution.
+3.1.4 Linear SSM with Time Shift State Transition Matrix
+The lower shift matrix is a structured matrix whose nonzero entries lie only on the subdiagonal. Let $\bm{A}_{d}$ in (2) denote such a matrix i.e. let
+$\bm{A}_{d}=\begin{bmatrix}0&0&\ldots&0&0\\ a_{1}&0&\ldots&0&0\\ 0&a_{2}&\ldots&0&0\\ \vdots&&\ddots&\vdots&\vdots\\ 0&0&\ldots&a_{n_{d}-1}&0\end{bmatrix}\text{ i.e., }[\bm{A}_{d}]_{i,j}=\begin{cases}a_{j}&\text{ for }i-1=j\\ 0&\text{ otherwise }\end{cases}$ (4a)
+where $\bm{A}_{d}]_{i,j}$ denotes the $(i,j)$-th entry of $\bm{A}_{d}$, and let $\bm{B}_{d}=[1,0,\ldots,0]^{T}$.
+Although the parameters $a_{i}$ are free in general, in this paper we set $a_{i}=1$ for all $i$.
+Under this choice, and following (3), $\bm{v}_{d}$ reduces to a simple temporal memory of the last $n$ inputs (a shift register), i.e.,
+$\bm{v}_{d}[t]=[i_{d}[t],i_{d}[t-1],i_{d}[t-2],\ldots,i_{d}[t-n_{d}+1]]^{T},$ (5)
+see Appendix 0.A.2 for a demonstration.
+3.2 Proposed Model
 Let $\bm{v}_{s}$ be the state variable of the general spiking neuron in (1) with parameter matrices $\bm{A}_{s},\bm{B}_{s},\bm{C}_{s}$ and state dimension $n_{s}$.
-Let $\bm{v}_{d}$ be the state variable of the time-shift SSM in Section IV-A2 with parameter matrices $\bm{A}_{d},\bm{B}_{d},\bm{C}_{d}$ and state dimension $n_{d}$.
+Let $\bm{v}_{d}$ be the state variable of the time-shift SSM in Section 3.1.4 with parameter matrices $\bm{A}_{d},\bm{B}_{d},\bm{C}_{d}$ and state dimension $n_{d}$.
 Our proposed spiking neuron model with state-delays is defined as:
@@
 $\bm{v}_{s}[t+1]$ $=\bm{A}_{s}\bm{v}_{s}[t]+\bm{B}_{s}i_{s}[t]+\bm{A}_{sd}\bm{v}_{d}[t]-\bm{R}s[t]$ (6b)
-$s[t]$ $=f_{\bf{\Theta}}(\bm{C}_{s}\bm{v}_{s}[t]+\bm{C}_{d}\bm{v}_{d}[t])$
-$=\begin{cases}1\text{ if }(\bm{C}_{s}\bm{v}_{s}[t]+\bm{C}_{d}\bm{v}_{d}[t])\in\bf{\Theta}\\ 0\text{ otherwise}\end{cases}$ (6c)
-Here, $\bm{A}_{sd}\in\mathbb{R}^{n_{s}\times n_{d}}$ gives the parameters that delayed inputs in the state $\bm{v}_{d}$ are multiplied with before they are added to the neuron state $\bm{v}_{s}$.
+$s[t]$ $=f_{\bf{\Theta}}(\bm{C}_{s}\bm{v}_{s}[t]+\bm{C}_{d}\bm{v}_{d}[t])=\begin{cases}1\text{ if }(\bm{C}_{s}\bm{v}_{s}[t]+\bm{C}_{d}\bm{v}_{d}[t])\in\bf{\Theta}\\ 0\text{ otherwise}\end{cases}$ (6c)
+Here, $\bm{A}_{sd}\in\mathbb{R}^{n_{s}\times n_{d}}$ denotes the parameters that delayed inputs in the state $\bm{v}_{d}$ are multiplied with before they are added to the neuron state $\bm{v}_{s}$.
 The above equations present a spiking neuron with additional access to temporal memory through the state variable $\bm{v}_{d}[t]$.
@@
 the number of directly accessible past quantities.
-The input $i_{s}$ to $\bm{v}_{s}$, as defined in Section IV-A1, represents the input spikes to the neuron.
+The input $i_{s}$ to $\bm{v}_{s}$, as defined in Section 3.1.1, represents the input spikes to the neuron.
 The input $i_{d}$ to $\bm{v}_{d}$ represents the input stored in the temporal memory.
-See Table I for an overview of the notation.
-Examples of the proposed models with LIF and adLIF neuron models are provided in Sections IV-B1 and IV-B2, respectively.
+See Table 1 for an overview of the notation.
+Examples of the proposed models with LIF, RLIF, adLIF and RadLIF neuron models are provided in Appendix 0.A.1.
 A high-level visualization of the proposed model is presented in Figure 2.
-TABLE I: Table of notation.
+Table 1: Table of notation.
 Notation Meaning
@@
 $h$ Number of neurons in a hidden layer
-Figure 2: High-level illustration of the proposed neuron model. The figure shows how the delay components are incorporated into the traditional SNN neuron model (LIF neuron shown).
-IV-B1 LIF and RLIF example
-The discrete-time Leaky Integrate-and-Fire (LIF) neuron, as defined in [bittar2022surrogate], with the additional proposed delays is given by:
+Figure 2: Proposed delay component incorporated into SNN neuron model (LIF).
+3.3 Delay parameters
+We now provide an overview of all parameters related to the additional delay variable $\bm{v}_{d}$ in (6).
+We have $\bm{A}_{d}$ and $\bm{B}_{d}$ as in Section 3.1.4, hence they are fixed and non-trainable.
+Under LIF/RLIF/adLIF/RadLIF neuron models (see Appendix 0.A.1), we do not use $\bm{C}_{d}$.
+Hence, the only additional (possibly trainable) parameter is $\bm{A}_{sd}$.
+In the rest of this section, we focus on $\bm{A}_{sd}$.
+The parameter matrix $\bm{A}_{sd}$ determines how the delay variable $\bm{v}_{d}$ contributes to the intrinsic neuron state $\bm{v}_{s}$ in (6). We consider four settings for its elements:
+  • $\bm{A}_{sd}=$ Ones, where $\bm{A}_{sd}]_{ij}=1$
+  • $\bm{A}_{sd}=$ Linear decay, where $\bm{A}_{sd}]_{ij}=(n_{d}-j)\frac{1}{n_{d}}$
+  • $\bm{A}_{sd}=$ Exponential decay, where $\bm{A}_{sd}]_{ij}=e^{-0.5j}$
+  • $\bm{A}_{sd}\sim$ Uniform distribution, where $\bm{A}_{sd}]_{ij}\sim U(0,1]$
+The equal sign == is used in the cases where the elements of $\bm{A}_{sd}$ are deterministic, while the sign $\sim$ is used in the case where the elements of $\bm{A}_{sd}$ are randomly drawn from a distribution.
+The matrix of delay parameters $\bm{A}_{sd}$ can be trainable or fixed.
+When it is fixed, its elements do not change during the training of the SNN and remain as initially set.
+When it is trainable, the initial set values of $\bm{A}_{sd}$ represent their initialization, and they change during the training phase without the constraint of preserving the distribution their initialization had, i.e., for instance if $\bm{A}_{sd}]_{ij}$ is initialized as ‘Ones’, the parameters do not stay all ones or all equal.
+In Section 5 we distinguish these cases by stating whether $\bm{A}_{sd}$ is non-trainable or trainable.
+Due to our parameter choices in Section 3.1.4, the delay variable $\bm{v}_{d}[t]$ is simply a vector containing the past $n_{d}$ scalar inputs $i_{d}[t]$, see (5).
+Consequently, the delay contribution $\bm{A}_{sd}\bm{v}_{d}[t]$ corresponds to a weighted combination of the past $n_{d}$ inputs $i_{d}[t]$ (see Figure 4 in Appendix 0.A.3 for illustration), i.e.,
+$\bm{A}_{sd}\bm{v}_{d}[t]\,=\,$ $[\bm{A}_{sd}]_{0}i_{d}[t]\,+\,[\bm{A}_{sd}]_{1}i_{d}[t-1]\,+\,\dots\,+\,[\bm{A}_{sd}]_{n_{d}-1}i_{d}[t-n_{d}+1].$ (7)
+Figure 3 illustrates how the coefficients $\bm{A}_{sd}]_{i}$ appear under the four settings described above.
+Figure 3: Visualization of the different parameter settings explored for the elements of $\bm{A}_{sd}$. For $\bm{A}_{sd}]_{ij}\sim$ Uniform, one random instance is shown as an example.
+3.4 Computational complexity
+The proposed neuron model with delays, given in (6), extends the baseline formulation in (1) by introducing an additional state variable $\bm{v}_{d}$ and its associated parameters $\bm{A}_{d}$, $\bm{B}_{d}$, $\bm{A}_{sd}$, and $\bm{C}_{d}$.
+The dimensionality of these matrices is determined by the delay order $n_{d}$.
+In this section, we analyze how this additional state and its parameters influence the computational complexity of the spiking neuron for the case of LIF and adLIF neuron as defined in Appendix 0.A.1.
+3.4.1 Number of trainable parameters
+The baseline definition of the LIF neuron model has $1$ trainable parameter, i.e., $\alpha$, while the adLIF neuron model has $4$ trainable parameters, i.e., $\alpha,\beta,a,b$.
+Let $l$ be the number of hidden layers, each with $h$ neurons, and let $c_{in}$ and $c_{out}$ be the number of input and output channels to the network.
+The trainable parameter counts for each network component are:
+  • feedforward synaptic connections: $c_{in}h+h^{2}(l-1)+hc_{out}$
+  • recurrent synaptic connections: $l(h^{2}-h)$
+  • neuron parameters: $lh$ for LIF and $4lh$ for adLIF
+  • normalization layers: $2(hl+c_{out})$
+  • delays: $n_{d}hl$ if $\bm{A}_{sd}$ trainable, 0 otherwise.
+From the above calculations, the delay mechanism introduces approx.
+$n_{d}hl$ additional parameters, but only if the matrix $\bm{A}_{sd}$ is trainable.
+If $\bm{A}_{sd}$ is non-trainable, the number of trainable parameters does not not increase i.e.
+it is equal to the number of trainable parameters of the neuron without delays.
+Even when $\bm{A}_{sd}$ is trainable, the number of delay-related parameters is small compared to the number of synaptic connections, which scale as $h^{2}$.
+Relatively small values of $n_{d}$ already achieve competitive performance (see Section 5).
+Hence, since $h$ is expected to be chosen as much larger than both $n_{d}$ and $l$ in LIF and adLIF networks, the additional computational and parametric overhead from incorporating delays is typically negligible.
+We now compare the number of trainable parameters introduced by our delay formulation with those reported in related work.
+In [7, Table 1], introducing delays increased the parameter count from 38.738.7k to 75.875.8k, i.e., by 37.137.1k parameters.
+For the same architecture, our approach increases the number of parameters by $n_{d}hl=n_{d}\times 128\times 2$.
+For $n_{d}=10$, this corresponds to about 2.62.6k parameters, an order of magnitude smaller than in [7].
+In [32, Section 2.1], each neuron learns a single delay shared across all outgoing connections, leading to $\mathcal{O}(h)$ parameter scaling.
+In our formulation, up to $n_{d}$ delay parameters can be learned per neuron while maintaining the same $\mathcal{O}(h)$ scaling, which is significantly more efficient than the $\mathcal{O}(h^{2})$ scaling of synaptic weights.
+3.4.2 Memory
+The delay state $\bm{v}_{d}\in\mathbb{R}^{n_{d}\times 1}$ acts as a memory buffer storing $n_{d}$ values per neuron.
+The intrinsic neuron state $\bm{v}_{s}\in\mathbb{R}^{n_{s}\times 1}$ already contains $n_{s}$ internal variables.
+For example, the LIF neuron has one state variable (membrane potential), $n_{s}=1$, while the adLIF includes an additional adaptation variable, giving $n_{s}=2$.
+Appending $\bm{v}_{d}$ to $\bm{v}_{s}$ in (6) increases the number of state variables per neuron from $n_{s}$ to $n_{s}+n_{d}$.
+At the network level, across $l$ hidden layers with $h$ neurons each, the total number of stored states grows from $n_{s}hl$ to $(n_{s}+n_{d})hl$.
+However, since $n_{s}$, $n_{d}$, and $l$ are typically much smaller than $h$ (see Section 3.4.1), the overall memory footprint in both cases scales primarily with $h$.
+4 Experimental Set-up
+4.0.1 Network Architecture and Training
+In all experiments, a network with two hidden layers is used with batch normalization layers and an accumulative output layer with cross entropy loss.
+Each model is trained using BPTT with surrogate gradient [27].
+We present further details for training and network parameters in Appendix 0.A.4.
+4.0.2 Dataset
+The Spiking Heidelberg Digits (SHD) dataset [5] consists of spike sequences derived from microphone recordings via the Lauscher cochlea model, covering $20$ classes (digits 0-99 in German and English) across $8156$ training and $2264$ test samples, each with $700$ input channels.
+The task is to classify the spoken digit from the input spike sequence.
+The Pre-processing details are in the Appendix 0.A.5.
+5 Results
+We discuss the numerical experiments in Tables 25, covering comparisons to literature (Section 5.1), the effect of delay order, delay parameters, training time, and neuron count (Sections 5.25.5), and the effect of trainable vs.
+non-trainable delays (Section 5.6).
+5.1 Comparison to literature
+In this section, we compare our best performance presented in this paper with the performances of SHD dataset with other SNN works which include delays, see Table 2.
+We observe that our model is comparable both in terms of performance and in terms of number of parameters used.
+Note that in general, the highest average accuracy reported over multiple initialization instances for the SHD dataset using SNNs is $95.07\%$ [17], which is included in Table 2.
+The benchmark performance using a non-spiking encoding of the same data and a recurrent neural network is $99.69\%$ [2].
+Table 2: Accuracy and number of parameters reported for delay-based SNN papers on the SHD dataset. M denotes millions of parameters.
+Paper Accuracy # params.
+Learning DCLS [17] 95.07 $\pm$ 0.24 % 0.2M
+Co-learning [7] 95.02 % 0.076M
+Learnable axonal delay [33] 93.55 % 0.21M
+Event-based Delay [25] 93.24 $\pm$ 1.00 % 0.5M (a)
+Temporal Hierarchy [26] 92.1 % (b) -
+DMP [34] 91.69 % 0.046M
+Heterogeneous Delays [32] 90.98 % 0.1M (c)
+Ours (adLIF, $n_{d}=5$, non-train. $\bm{A}_{sd}$) 94.1 $\pm$ 0.5 % 0.04M
+  • (a)
+    Read from [25, Fig. 5];
+  • (b)
+    Read from [26, Fig. 5];
+  • (c)
+    Read from [32, Fig. 1c].
+5.2 Order of delays
+In this section, we discuss the varying order of delays i.e., we compare $n_{d}=0$, $5$ and $10$, see Table 3.
+We observe that increasing the delay order from $n_{d}=0$ (i.e.
+no delay, hence the same as the traditional neuron types) to $n_{d}=5$ or $n_{d}=10$ either maintains or improves performance for LIF, adLIF, and RadLIF neurons.
+For example for $\bm{A}_{sd}$ non-trainable, the adLIF baseline accuracy with $n_{d}=0$ is $92.0\pm 0.4\%$, while for $\bm{A}_{sd}=\text{Uniform}$ and $n_{d}=5$ the accuracy rises to $94.1\pm 0.5\%$, indicating an improvement beyond the standard deviation range.
+In contrast, RLIF neuron either maintains its performance or shows only a slight decrease when delays are introduced.
+For example, for $\bm{A}_{sd}$ non-trainable, RLIF baseline accuracy with $n_{d}=0$ is $90.6\pm 0.5\%$, while for $\bm{A}_{sd}=\text{Ones}$ and $n_{d}=5$ the accuracy drops to $87.7\pm 1.1\%$, indicating a performance drop beyond the standard deviation range.
+When comparing $n_{d}=5$ and $n_{d}=10$, the results remain within the standard deviation of each other, suggesting that increasing the delay order to $n_{d}=10$ does not yield additional benefits beyond what $n_{d}=5$ already captures.
+Table 3: Test accuracy on the SHD dataset across combinations of neuron model, delay parameters $\bm{A}_{sd}$, and delay order $n_{d}$.
+All networks consist of two hidden layers with $h=128$ neurons each.
+Bold indicates the best result per row.
+(a) $\bm{A}_{sd}$ is non-trainable.
+Neuron Model No delay $\bm{A}_{sd}=$ Ones $\bm{A}_{sd}=$ Lin. Decay $\bm{A}_{sd}=$ Exp. Decay $\bm{A}_{sd}\sim$ Uniform
+$n_{d}=5$ $n_{d}=10$ $n_{d}=5$ $n_{d}=10$ $n_{d}=5$ $n_{d}=10$ $n_{d}=5$ $n_{d}=10$
+LIF 85.9 $\pm$ 1.2 86.6 $\pm$ 1.2 83.8 $\pm$ 2.0 86.6 $\pm$ 0.8 85.9 $\pm$ 0.8 86.7 $\pm$ 0.7 86.0 $\pm$ 0.9 87.2 $\pm$ 1.6 86.7 $\pm$ 0.8
+RLIF 90.6 $\pm$ 0.5 87.7 $\pm$ 1.1 87.1 $\pm$ 1.1 88.8 $\pm$ 1.0 87.8 $\pm$ 0.6 87.9 $\pm$ 1.7 89.0 $\pm$ 0.7 89.6 $\pm$ 1.0 88.0 $\pm$ 0.9
+adLIF 92.0 $\pm$ 0.4 92.1 $\pm$ 0.2 91.6 $\pm$ 0.4 93.3 $\pm$ 0.9 91.8 $\pm$ 0.8 93.3 $\pm$ 0.4 93.0 $\pm$ 0.9 94.1 $\pm$ 0.5 92.1 $\pm$ 1.3
+RadLIF 93.4 $\pm$ 1.1 91.6 $\pm$ 2.6 91.2 $\pm$ 2.2 92.8 $\pm$ 0.7 92.4 $\pm$ 0.9 93.8 $\pm$ 0.6 93.5 $\pm$ 0.4 93.1 $\pm$ 1.2 92.1 $\pm$ 1.2
+(b) $\bm{A}_{sd}$ is trainable.
+Neuron Model $\bm{A}_{sd}=$ Ones init. $\bm{A}_{sd}=$ Lin. Decay init. $\bm{A}_{sd}=$ Exp. Decay init. $\bm{A}_{sd}\sim$ Uniform init.
+$n_{d}=5$ $n_{d}=10$ $n_{d}=5$ $n_{d}=10$ $n_{d}=5$ $n_{d}=10$ $n_{d}=5$ $n_{d}=10$
+LIF 87.7 $\pm$ 0.9 88.2 $\pm$ 1.0 88.4 $\pm$ 1.2 89.1 $\pm$ 0.3 89.1 $\pm$ 0.8 89.6 $\pm$ 1.1 88.4 $\pm$ 1.5 89.5 $\pm$ 0.5
+RLIF 88.5 $\pm$ 1.7 87.9 $\pm$ 1.3 89.0 $\pm$ 0.5 89.5 $\pm$ 1.2 89.1 $\pm$ 0.6 90.3 $\pm$ 0.9 88.9 $\pm$ 0.5 89.7 $\pm$ 0.8
+adLIF 92.7 $\pm$ 0.7 91.9 $\pm$ 0.5 93.1 $\pm$ 0.4 92.7 $\pm$ 0.5 93.6 $\pm$ 0.2 93.6 $\pm$ 1.1 93.3 $\pm$ 0.6 92.6 $\pm$ 0.5
+RadLIF 93.4 $\pm$ 0.8 92.4 $\pm$ 1.3 94.0 $\pm$ 0.2 93.3 $\pm$ 0.6 93.4 $\pm$ 0.5 93.7 $\pm$ 0.6 93.7 $\pm$ 0.6 92.3 $\pm$ 1.0
+5.3 Delay parameters
+In this section we investigate the impact of using various delay parameter settings for $\bm{A}_{sd}$.
+In particular, we explore $\bm{A}_{sd}=$ Ones, Linear Decay, Exponential Decay, and $\bm{A}_{sd}\sim$ Uniform, see Section 3.3 for more details of the setting.
+Table 3 provides the associated results.
+No consistent trend for delay weight parametrization is observed across all $n_{d}$ values and neuron types, with most accuracy values falling within one standard deviation of each other and remaining differences appearing model-specific.
+For instance, the adLIF neuron achieves its best non-trainable $\bm{A}_{sd}$ performance with the Uniform setting, while the Exponential Decay initialization yields the highest accuracy in the trainable case.
+5.4 Training time
+In this section, we discuss the additional training time introduced by using delays on NVIDIA Tesla T4 GPU with 16GB RAM, see Table 4.
+We first focus on the increase in training time when using $n_{d}=5$ compared to the baseline with no delay.
+The smallest ratio increase, $\frac{13.1}{8.6}=1.5$, occurs for the adLIF model with a non-trainable $\bm{A}_{sd}$, while the largest, $\frac{12.5}{6.5}=1.8$, is observed for the LIF model with a trainable $\bm{A}_{sd}$.
+Increasing $n_{d}$ from $5$ to $10$ results in only a negligible change in training time.
+Table 4: Training time in minutes as average over the various delay parameters $\bm{A}_{sd}$ settings for fixed $n_{d}$ from Table 3.
+$\bm{A}_{sd}$ non-trainable $\bm{A}_{sd}$ trainable
+Neuron Model No delay $n_{d}=5$ $n_{d}=10$ $n_{d}=5$ $n_{d}=10$
+LIF 6.5 $\pm$ 0.1 11.7 $\pm$ 0.1 12.0 $\pm$ 0.2 12.5 $\pm$ 0.2 12.7 $\pm$ 0.3
+RLIF 7.9 $\pm$ 0.1 13.1 $\pm$ 0.2 12.9 $\pm$ 0.1 13.9 $\pm$ 0.2 13.7 $\pm$ 0.1
+adLIF 8.6 $\pm$ 0.1 13.1 $\pm$ 0.3 13.4 $\pm$ 0.6 14.1 $\pm$ 0.2 13.9 $\pm$ 0.1
+RadLIF 9.6 $\pm$ 0.1 15.4 $\pm$ 0.2 15.1 $\pm$ 0.1 15.9 $\pm$ 0.4 16.0 $\pm$ 0.3
+5.5 Number of neurons vs order of delays
+In this section, we discuss the combined effect of the number of neurons in the hidden layers $h$ and the order of delays $n_{d}$, see Table 5.
+The impact of increasing $n_{d}$ on the performance depends on the number of neurons $h$.
+For $h\geq 32$, using delays ($n_{d}=10,50$ or $100$) maintains performance at a level similar to no-delay or slightly reduces it.
+In contrast, for smaller networks with $h\leq 16$, using delays yields a clear improvement over the baseline no-delay performance.
+Note that in our implementation, the input sequences of the SHD dataset have a length of $100$.
+Consequently, $n_{d}=100$ allows the network to access the entire temporal history of each sample at every time step.
+Table 5: Test accuracy on the SHD dataset across combinations of neuron count $h$ and delay order $n_{d}$ for adLIF neurons with $\bm{A}_{sd}\sim\text{Uniform}$.
+Bold indicates the best result per row.
+$\bm{A}_{sd}$ non-trainable $\bm{A}_{sd}$ trainable
+$h$ No delay $n_{d}=10$ $n_{d}=50$ $n_{d}=100$ $n_{d}=10$ $n_{d}=50$ $n_{d}=100$
+512 92.5 $\pm$ 0.7 92.7 $\pm$ 0.9 90.8 $\pm$ 1.5 90.4 $\pm$ 1.0 93.6 $\pm$ 0.3 92.2 $\pm$ 0.8 92.7 $\pm$ 0.9
+256 93.1 $\pm$ 1.1 92.1 $\pm$ 0.5 90.6 $\pm$ 1.3 89.9 $\pm$ 0.6 93.4 $\pm$ 0.6 92.2 $\pm$ 0.7 91.7 $\pm$ 1.2
+128 92.0 $\pm$ 0.4 92.1 $\pm$ 1.3 89.8 $\pm$ 1.2 89.4 $\pm$ 0.6 92.6 $\pm$ 0.5 92.3 $\pm$ 0.5 91.9 $\pm$ 1.0
+64 91.3 $\pm$ 1.1 91.8 $\pm$ 1.3 89.6 $\pm$ 1.1 89.4 $\pm$ 1.8 92.7 $\pm$ 1.1 91.9 $\pm$ 1.0 91.5 $\pm$ 1.0
+32 88.7 $\pm$ 2.2 90.2 $\pm$ 1.3 86.8 $\pm$ 1.4 88.2 $\pm$ 1.2 90.5 $\pm$ 1.3 90.5 $\pm$ 0.8 90.4 $\pm$ 0.9
+16 78.1 $\pm$ 1.0 86.2 $\pm$ 1.0 84.1 $\pm$ 1.2 83.0 $\pm$ 1.0 86.9 $\pm$ 1.2 86.6 $\pm$ 0.7 87.1 $\pm$ 1.2
+8 60.1 $\pm$ 7.4 72.6 $\pm$ 2.5 70.1 $\pm$ 2.6 73.8 $\pm$ 1.9 75.3 $\pm$ 1.6 77.3 $\pm$ 2.2 80.3 $\pm$ 1.3
+5.6 Trainable vs non-trainable delay parameters
+In this section, we compare results with trainable and non-trainable $\bm{A}_{sd}$ matrices.
+Tables 3(a) and 3(b) show that the accuracy values are generally within one standard deviation of each other, except for the LIF case where trainable $\bm{A}_{sd}$ provides a clear benefit.
+Similarly, Table 5 shows no meaningful difference for $h\geq 32$.
+In contrast, for smaller networks ($h\leq 16$), the difference can be significant, for example, at $h=8$ and $n_{d}=100$, trainable $\bm{A}_{sd}$ improves accuracy from $73.8\%$ to $80.3\%$.
+Table 4 shows that the additional training time for trainable $\bm{A}_{sd}$ is negligible (<<1 min), further supporting the low computational overhead discussed in Section 3.4.1.
+6 Discussions
+The limited state dimensionality of LIF and adLIF neurons may restrict their ability to fully exploit the proposed delay mechanism; richer neuron models with more internal dynamics could deliver further benefits.
+No dedicated hyperparameter optimization (HPO) was performed, which could further improve performance and alter observations.
+Finally, whether the proposed delay neuron models can be efficiently implemented on neuromorphic hardware, as has been explored for some mechanisms [14], remains an open question.
+7 Conclusion
+In this work, we propose a general framework for incorporating delays into spiking neuron dynamics by introducing additional state variables that encode a finite temporal history, enabling seamless integration with a wide range of neuron models.
+We analyze the effect of delay order, trainable versus fixed delay parameters, and the trade-off between delay order and network size.
+The results show that the proposed mechanism achieves performance comparable to existing delay-based approaches while introducing minimal additional trainable parameters or not at all.
+{credits}
+7.0.1 Acknowledgements
+S.
+Karilanova and A.
+Özçelikkale acknowledge the support of Center for Interdisciplinary Mathematics (CIM), Uppsala University.
+The computations were enabled by resources provided by the National Academic Infrastructure for Supercomputing in Sweden (NAISS), partially funded by the Swedish Research Council through grant agreement no.
+2022-06725.
+Appendix 0.A Appendix
+0.A.1 Proposed Method applied to LIF/AdLIF
+0.A.1.1 LIF and RLIF example
+The discrete-time Leaky Integrate-and-Fire (LIF) neuron, as defined in [1], with the additional proposed delays is given by:
 $\bm{v}_{d}[t+1]$ $=\bm{A}_{d}\bm{v}_{d}[t]+\bm{B}_{d}i_{d}[t]$ (8a)
@@
 $s[t]$ $=f_{\theta}(u[t])=\begin{cases}1\text{ if }u[t]\geq\theta\\ 0\text{ otherwise.}\end{cases}$ (8c)
-where $i_{s}[t]=\bm{W}\times\bm{s}^{l-1}[t]+\bm{V}\times\bm{s}^{l}[t]$ for RLIF and $i_{s}[t]=\bm{W}\times\bm{s}^{l-1}[t]$ for LIF.
+where $i_{s}[t]=\bm{W}\times\bm{s}^{l-1}[t]+\bm{V}\times\bm{s}^{l}[t]$ for Recurrent LIF (RLIF) and $i_{s}[t]=\bm{W}\times\bm{s}^{l-1}[t]$ for LIF.
 For both RLIF and LIF, we have $i_{d}[t]=\bm{W}\times\bm{s}^{l-1}[t]$.
-IV-B2 adLIF and RadLIF example
-The Adaptive Leaky Integrate-and-Fire (adLIF) neuron is an extension of the LIF neuron with an additional recovery state variable, $w[t]$ [bittar2022surrogate].
+0.A.1.2 adLIF and RadLIF example
+The Adaptive Leaky Integrate-and-Fire (adLIF) neuron is an extension of the LIF neuron with an additional recovery state variable, $w[t]$ [1].
 Discrete-time adLIF neuron with the additional proposed delays is:
@@
 $s[t]=$ $f_{\theta}(u[t])=\begin{cases}1\text{ if }u[t]\geq\theta\\ 0\text{ otherwise.}\end{cases}$ (9d)
-The inputs $i_{s}[t]$ and $i_{d}[t]$ are as defined for RLIF in Section IV-B1 with a correspondence between LIF and adLIF, and RLIF and RadLIF.
-IV-C Delay parameters
-We now provide an overview of all parameters related to the additional delay variable $\bm{v}_{d}$ in (6).
-We have $\bm{A}_{d}$ given by (3) under $a_{i}=1$ and $\bm{B}_{d}$ given by (4), hence they are fixed and non-trainable.
-Under LIF/RLIF neuron models of (7) and adLIF/RadLIF neuron models of (8), we do not use $\bm{C}_{d}$.
-Hence, the only additional (possibly trainable) parameter is $\bm{A}_{sd}$.
-In the rest of this section, we focus on $\bm{A}_{sd}$.
-The matrix of parameters $\bm{A}_{sd}$ plays a central role in our development, as it determines the contribution of the delay variable $\bm{v}_{d}$ when added to the intrinsic neuron state $\bm{v}_{s}$, see (6).
-We consider four different settings for the elements of $\bm{A}_{sd}$:
-  1. 1.
-    $\bm{A}_{sd}=$ Ones, where $\bm{A}_{sd}]_{ij}=1$
-  2. 2.
-    $\bm{A}_{sd}=$ Linear decay, where $\bm{A}_{sd}]_{ij}=(n_{d}-j)\frac{1}{n_{d}}$
-  3. 3.
-    $\bm{A}_{sd}=$ Exponential decay, where $\bm{A}_{sd}]_{ij}=e^{-0.5j}$
-  4. 4.
-    $\bm{A}_{sd}\sim$ Uniform distribution, where $\bm{A}_{sd}]_{ij}\sim U(0,1]$
-The equal sign == is used in the cases where the elements of $\bm{A}_{sd}$ are deterministic, while the sign $\sim$ is used in the case where the elements of $\bm{A}_{sd}$ are random.
-The matrix of delay parameters $\bm{A}_{sd}$ can be trainable or fixed.
-When it is fixed, its elements do not change during the training of the SNN and remain as initially set.
-When it is trainable, the initial set values of $\bm{A}_{sd}$ represent their initialization, and they change during the training phase without the constraint of preserving the distribution their initialization had, i.e., for instance if $\bm{A}_{sd}]_{ij}$ is initialized as ’Ones’, the parameters do not stay all ones or all equal.
-In Section VI we distinguish these cases by stating whether $\bm{A}_{sd}$ is non-trainable or trainable.
-We now illustrate how the matrix $\bm{A}_{sd}$ determines the contribution of the delayed component $\bm{v}_{d}[t]$ in (5).
-The contributing delay component $\bm{A}_{sd}\bm{v}_{d}[t]$ is simply a weighted average of the past $n_{d}$ inputs $i_{d}[t]$.
-In particular, we have the following cases:
-  1. 1.
-    For $\bm{A}_{sd}=$ Ones:
-    $\bm{A}_{sd}\bm{v}_{d}[t]=i_{d}[t]+i_{d}[t-1]+i_{d}[t-2]+$...
-  2. 2.
-    For $\bm{A}_{sd}=$ Linear Decay: For instance, with $n_{d}=5$:
-    $\bm{A}_{sd}\bm{v}_{d}[t]=i_{d}[t]+0.8i_{d}[t-1]+0.6i_{d}[t-2]+$...
-  3. 3.
-    For $\bm{A}_{sd}=$ Exponential Decay: For instance, with $n_{d}=5$:
-    $\bm{A}_{sd}\bm{v}_{d}[t]=i_{d}[t]+0.60i_{d}[t-1]+0.34i_{d}[t-2]+$...
-  4. 4.
-    For $\bm{A}_{sd}\sim$ Uniform. Here, the elements are random. A particular realization could be:
-    $\bm{A}_{sd}\bm{v}_{d}[t]=0.1i_{d}[t]+0.9i_{d}[t-1]+0.2i_{d}[t-2]+$...
-IV-D Computational complexity
-The proposed neuron model with delays, given in (6), extends the baseline formulation in (1) by introducing an additional state variable $\bm{v}_{d}$ and its associated parameters $\bm{A}_{d}$, $\bm{B}_{d}$, $\bm{A}_{sd}$, and $\bm{C}_{d}$.
-The dimensionality of these matrices is determined by the delay order $n_{d}$.
-In this section, we analyze how this additional state and its parameters influence the computational complexity of the spiking neuron for the case of LIF and adLIF neuron as defined in Section IV-B1 and IV-B2.
-IV-D1 Number of trainable parameters
-The baseline definition of the LIF neuron model has $1$ trainable parameter, i.e., $\alpha$, while the adLIF neuron model has $4$ trainable parameters, i.e., $\alpha,\beta,a,b$.
-Let $l$ be the number of hidden layers, each with $h$ number of neurons.
-Let $c_{in}$ and $c_{out}$ be the number of input and output channels to the network.
-Then, the number of trainable parameters introduced by individual components of the networks is given by:
-  • feedforward synaptic connections: $c_{in}h+h^{2}(l-1)+hc_{out}$
-  • recurrent synaptic connections: $l(h^{2}-h)$
-  • neuron parameters: $lh$ for LIF and =4lh=4lh for adLIF
-  • normalization layers: $2(hl+c_{out})$
-  • delays: $n_{d}hl$ if $\bm{A}_{sd}$ trainable, 0 otherwise.
-From the above calculations, we observe that the number of additional parameters introduced by the delay mechanism is  $n_{d}hl$, and only if the matrix $\bm{A}_{sd}$ is trainable.
-Even in this case, the number of delay-related parameters remains relatively small compared to the number of synaptic connections, which scale as $h^{2}$.
-Relatively small values of $n_{d}$ get competitive performance, see Section.
-VI.
-Hence, since $h$ is much larger than both  $n_{d}$ or $l$ in both LIF and adLIF networks, the added computational and parametric overhead of incorporating delays is negligible.
-In the remainder of this section, we compare the number of trainable parameters introduced by our proposed delay modeling with those reported in related work.
-In [deckers2024co, Table 1], the inclusion of delays increased the total number of parameters from 38.738.7k to 75.875.8k i.e., by 37.137.1k parameters.
-For the same network architecture we increase the total number of parameters by $n_{d}hl=n_{d}\times 128\times 2$ parameters.
-For $n_{d}=10$, this corresponds to approximately 2.62.6k parameters, which is an order of magnitude smaller than in [deckers2024co].
-[sun2025exploitingheterogeneousdelaysefficient, Section 2.1] reports that each neuron maintains a single learned delay shared across all outgoing connections, resulting in $\mathcal{O}(h)$ parameter scaling.
-In our formulation, up to $n_{d}$ delay-related parameters can be trained per neuron, where the overall scaling remains $\mathcal{O}(h)$.
-As noted in [sun2025exploitingheterogeneousdelaysefficient], this is considerably more efficient than the $\mathcal{O}(h^{2})$ scaling of synaptic weights.
-IV-D2 Memory
-The state variable $\bm{v}_{d}\in\mathbb{R}^{n_{d}\times 1}$ can be seen as a memory buffer that stores $n_{d}$ values per neuron.
-The intrinsic neuron state $\bm{v}_{s}\in\mathbb{R}^{n_{s}\times 1}$ already contains $n_{s}$ internal variables that evolve over time.
-For example, the standard LIF neuron has a single state variable, the membrane potential, resulting in $n_{s}=1$.
-The adLIF, compared to the LIF, includes an additional threshold–adaptation variable, resulting in $n_{s}=2$.
-When the delay variable $\bm{v}_{d}$ is appended to the intrinsic state $\bm{v}_{s}$ in (6), the number of state variables for each neuron increases from $n_{s}$ to $n_{s}+n_{d}$.
-At the network level, this increase applies to all hidden layers $l$ with $h$ neurons, increasing the total number of stored state variables from $n_{s}hl$ to $(n_{s}+n_{d})hl$.
-However, as noted in Section IV-D1, the values $n_{s},n_{d}$ and $l$ are all much smaller than $h$.
-Consequently, in both cases, with or without delays, the memory footprint scales predominantly with $h$.
-V Experimental Set-up
-V-1 Network Architecture and Training
-In all experiments, a network with two hidden layers is used with batch normalization layers [ioffe2015batchnormalizationacceleratingdeep] and an accumulative output layer with cross entropy loss.
-Each model is trained using BPTT with surrogate gradient [neftci2019surrogate], AdamW optimizer [loshchilov2019decoupledweightdecayregularization] and cosine learning rate scheduler.
-We present further details for training and network parameters in Section IX-A.
-V-2 Dataset
-The Spiking Heidelberg Digits (SHD) dataset was introduced to standardize the evaluation of neuromorphic learning models [shd_ssc_dataset].
-It consists of spike sequences derived from microphone recordings using the Lauscher artificial cochlea model.
-The dataset contains $20$ classes of spoken digits 0 to $9$ in both German and English.
-It has $8156$ training and $2264$ test samples.
-Each sample is represented by $700$ input channels.
-This dataset is primarily used for classification tasks, where the goal is to identify the spoken digit from the input spike sequence.
-Further details on dataset preprocessing and usage are provided in Section IX-B.
-VI Results
-In this section, we discuss the numerical experiments presented in Tables II, III, IV, V.
-Specifically in Section VI-A we comment on the performance of our proposed method in comparison to the works in the literature.
-In Sections VI-B, VI-C, VI-D, and VI-E we comment on the varying order of delays, the delay parameters, the corresponding network training time, and the varying number of neurons used, respectively.
-In the final Section VI-F we discuss the effect of trainable and non-trainable delay parameters.
-TABLE II: Accuracy and number of parameters reported for delay-based SNN papers on the SHD dataset. M denotes millions of parameters.
-Paper Accuracy # params.
-Learning DCLS [hammouamri2023learning] 95.07 $\pm$ 0.24 % 0.2M
-Co-learning [deckers2024co] 95.02 % 0.076M
-Learnable axonal delay [10.3389/fnins.2023.1275944] 93.55 % 0.21M
-Event-based Delay [meszaros2025efficient] 93.24 $\pm$ 1.00 % 0.5M (a)
-Temporal Hierarchy [moro2024roletemporalhierarchyspiking] 92.1 % (b) -
-Heterogeneous Delays [sun2025exploitingheterogeneousdelaysefficient] 90.98 % 0.1M (c)
-Our (adLIF, $n_{d}=5$, non-train. $\bm{A}_{sd}$) 94.1 $\pm$ 0.5 % 0.04M
-  • (a)
-    Read from [meszaros2025efficient, Fig. 5]
-  • (b)
-    Read from [moro2024roletemporalhierarchyspiking, Fig. 5]
-  • (c)
-    Read from [Fig.1c] [sun2025exploitingheterogeneousdelaysefficient]
-TABLE III: Test accuracy on the SHD dataset across combinations of neuron model, delay parameters $\bm{A}_{sd}$, and delay order $n_{d}$.
-All networks consist of two hidden layers with $h=128$ neurons each.
-Bold values indicate the best performance for each neuron model (per row).
-(a) $\bm{A}_{sd}$ is non-trainable.
-Neuron Model No delay $\bm{A}_{sd}=$ Ones $\bm{A}_{sd}=$ Lin. Decay $\bm{A}_{sd}=$ Exp. Decay $\bm{A}_{sd}\sim$ Uniform
-$n_{d}=5$ $n_{d}=10$ $n_{d}=5$ $n_{d}=10$ $n_{d}=5$ $n_{d}=10$ $n_{d}=5$ $n_{d}=10$
-LIF 85.9 $\pm$ 1.2 % 86.6 $\pm$ 1.2 % 83.8 $\pm$ 2.0 % 86.6 $\pm$ 0.8 % 85.9 $\pm$ 0.8 % 86.7 $\pm$ 0.7 % 86.0 $\pm$ 0.9 % 87.2 $\pm$ 1.6 % 86.7 $\pm$ 0.8 %
-RLIF 90.6 $\pm$ 0.5 % 87.7 $\pm$ 1.1 % 87.1 $\pm$ 1.1 % 88.8 $\pm$ 1.0 % 87.8 $\pm$ 0.6 % 87.9 $\pm$ 1.7 % 89.0 $\pm$ 0.7 % 89.6 $\pm$ 1.0 % 88.0 $\pm$ 0.9 %
-adLIF 92.0 $\pm$ 0.4 % 92.1 $\pm$ 0.2 % 91.6 $\pm$ 0.4 % 93.3 $\pm$ 0.9 % 91.8 $\pm$ 0.8 % 93.3 $\pm$ 0.4 % 93.0 $\pm$ 0.9 % 94.1 $\pm$ 0.5 % 92.1 $\pm$ 1.3 %
-RadLIF 93.4 $\pm$ 1.1 % 91.6 $\pm$ 2.6 % 91.2 $\pm$ 2.2 % 92.8 $\pm$ 0.7 % 92.4 $\pm$ 0.9 % 93.8 $\pm$ 0.6 % 93.5 $\pm$ 0.4 % 93.1 $\pm$ 1.2 % 92.1 $\pm$ 1.2 %
-(b) $\bm{A}_{sd}$ is trainable.
-Neuron Model $\bm{A}_{sd}=$ Ones init. $\bm{A}_{sd}=$ Lin. Decay init. $\bm{A}_{sd}=$ Exp. Decay init. $\bm{A}_{sd}\sim$ Uniform init.
-$n_{d}=5$ $n_{d}=10$ $n_{d}=5$ $n_{d}=10$ $n_{d}=5$ $n_{d}=10$ $n_{d}=5$ $n_{d}=10$
-LIF 87.7 $\pm$ 0.9 % 88.2 $\pm$ 1.0 % 88.4 $\pm$ 1.2 % 89.1 $\pm$ 0.3 % 89.1 $\pm$ 0.8 % 89.6 $\pm$ 1.1 % 88.4 $\pm$ 1.5 % 89.5 $\pm$ 0.5 %
-RLIF 88.5 $\pm$ 1.7 % 87.9 $\pm$ 1.3 % 89.0 $\pm$ 0.5 % 89.5 $\pm$ 1.2 % 89.1 $\pm$ 0.6 % 90.3 $\pm$ 0.9 % 88.9 $\pm$ 0.5 % 89.7 $\pm$ 0.8 %
-adLIF 92.7 $\pm$ 0.7 % 91.9 $\pm$ 0.5 % 93.1 $\pm$ 0.4 % 92.7 $\pm$ 0.5 % 93.6 $\pm$ 0.2 % 93.6 $\pm$ 1.1 % 93.3 $\pm$ 0.6 % 92.6 $\pm$ 0.5 %
-RadLIF 93.4 $\pm$ 0.8 % 92.4 $\pm$ 1.3 % 94.0 $\pm$ 0.2 % 93.3 $\pm$ 0.6 % 93.4 $\pm$ 0.5 % 93.7 $\pm$ 0.6 % 93.7 $\pm$ 0.6 % 92.3 $\pm$ 1.0 %
-VI-A Comparison to literature
-In this section, we compare our best performance presented in this paper with the performances of SHD dataset with other SNN works which include delays, see Table II.
-We observe that our model is comparable both in terms of performance and in terms of number of parameters used.
-Note that in general, the highest average accuracy reported over multiple initialization instances for the SHD dataset using SNNs is $95.07\%$ [hammouamri2023learning], which is included in Table II.
-The benchmark performance using a non-spiking encoding of the same data and a recurrent neural network is $99.69\%$ [boeshertz2024accuratemappingrnnsneuromorphic].
-VI-B Order of delays
-In this section, we discuss the varying order of delays i.e., we compare $n_{d}=0$, $5$ and $10$, see Table III.
-We observe that increasing the delay order from $n_{d}=0$ (i.e.
-no delay, hence the same as the traditional neuron types) to $n_{d}=5$ or $n_{d}=10$ in the LIF, adLIF, and RadLIF neurons either maintains (i.e.
-the performance with and without delay are within standard deviations of each other) or improves performance.
-For instance, in Table III(a) for the adLIF neuron model, the baseline accuracy with $n_{d}=0$ is $92.0\pm 0.4\%$, while for $\bm{A}_{sd}=\text{Uniform}$ and $n_{d}=5$ the accuracy rises to $94.1\pm 0.5\%$, indicating an improvement beyond the standard deviation range.
-In contrast, the RLIF neuron either maintains its performance or shows only a slight decrease when delays are introduced.
-For example, in Table III(a) for the RLIF neuron model, the baseline accuracy with $n_{d}=0$ is $90.6\pm 0.5\%$, while for $\bm{A}_{sd}=\text{Ones}$ and $n_{d}=5$ the accuracy drops to $87.7\pm 1.1\%$, indicating a performance drop beyond the standard deviation range.
-When comparing $n_{d}=5$ and $n_{d}=10$, the results remain within the standard deviation of each other, suggesting that increasing the delay order to $n_{d}=10$ does not yield additional benefits beyond what $n_{d}=5$ already captures.
-VI-C Delay parameters
-In this section we investigate the impact of using various delay parameter settings for $\bm{A}_{sd}$.
-In particular, we explore $\bm{A}_{sd}=$ Ones, Linear Decay, Exponential Decay, and $\bm{A}_{sd}\sim$ Uniform, see Section IV-C for more details of the setting.
-Table III provides the associated results.
-We do not observe a consistent trend that holds in all values of $n_{d}$ and in all neuron types.
-Most of the accuracy values are within the standard deviation of each other.
-The remaining differences appear to be model-specific.
-For instance, considering the average performance of the adLIF neuron, the best configuration for a non-trainable $\bm{A}_{sd}$ is the Uniform setting, whereas for a trainable $\bm{A}_{sd}$, the Exponential Decay initialization yields the highest accuracy.
-VI-D Training time
-In this section, we discuss the additional training time introduced by using delays, see Table IV.
-We first focus on the increase in training time when using $n_{d}=5$ compared to the baseline with no delay.
-The smallest ratio increase, $\frac{13.1}{8.6}=1.5$, occurs for the adLIF model with a non-trainable $\bm{A}_{sd}$, while the largest, $\frac{12.5}{6.5}=1.8$, is observed for the LIF model with a trainable $\bm{A}_{sd}$.
-Increasing $n_{d}$ from $5$ to $10$ results in only a negligible change in training time.
-TABLE IV: Training time in minutes as average over the various delay parameters $\bm{A}_{sd}$ settings for fixed $n_{d}$ from Table III.
-(a) $\bm{A}_{sd}$ non-trainable
-Neuron Model No delay $n_{d}=5$ $n_{d}=10$
-LIF 6.5 $\pm$ 0.1 min. 11.7 $\pm$ 0.1 min. 12.0 $\pm$ 0.2 min.
-RLIF 7.9 $\pm$ 0.1 min. 13.1 $\pm$ 0.2 min. 12.9 $\pm$ 0.1 min.
-adLIF 8.6 $\pm$ 0.1 min. 13.1 $\pm$ 0.3 min. 13.4 $\pm$ 0.6 min.
-RadLIF 9.6 $\pm$ 0.1 min. 15.4 $\pm$ 0.2 min. 15.1 $\pm$ 0.1 min.
-(b) $\bm{A}_{sd}$ trainable
-Neuron Model $n_{d}=5$ $n_{d}=10$
-LIF 12.5 $\pm$ 0.2 min. 12.7 $\pm$ 0.3 min.
-RLIF 13.9 $\pm$ 0.2 min. 13.7 $\pm$ 0.1 min.
-adLIF 14.1 $\pm$ 0.2 min. 13.9 $\pm$ 0.1 min.
-RadLIF 15.9 $\pm$ 0.4 min. 16.0 $\pm$ 0.3 min.
-VI-E Number of neurons vs order of delays
-In this section, we discuss the combined effect of the number of neurons in the hidden layers $h$ and the order of delays $n_{d}$, see Table V.
-The impact of increasing $n_{d}$ on the performance depends on the number of neurons $h$.
-For $h\geq 32$, using delays ($n_{d}=10,50$ or $100$) maintains performance at a level similar to no-delay or slightly reduces it.
-In contrast, for smaller networks with $h\leq 16$, using delays yields a clear improvement over the baseline no-delay performance.
-It is worth noting that, in our implementation, the input sequences of the SHD dataset have a length of $100$.
-Consequently, setting $n_{d}=100$ allows the network to access the entire temporal history of each sample at every time step.
-TABLE V: Test accuracy on the SHD dataset across combinations of number of neurons $h$ and delay order $n_{d}$.
-All networks consist of two hidden layers with $h$ neurons each, adLIF neuron model and $\bm{A}_{sd}\sim$ Uniform used as a fixed value in the $\bm{A}_{sd}$ non-trainable case and as initialization in the $\bm{A}_{sd}$ trainable case.
-Bold values indicate the best performance for each $h$ value (per row).
-(a) $\bm{A}_{sd}$ non-trainable.
-$h$ No delay $n_{d}=10$ $n_{d}=50$ $n_{d}=100$
-512 92.5 $\pm$ 0.7 % 92.7 $\pm$ 0.9 % 90.8 $\pm$ 1.5 % 90.4 $\pm$ 1.0 %
-256 93.1 $\pm$ 1.1 % 92.1 $\pm$ 0.5 % 90.6 $\pm$ 1.3 % 89.9 $\pm$ 0.6 %
-128 92.0 $\pm$ 0.4 % 92.1 $\pm$ 1.3 % 89.8 $\pm$ 1.2 % 89.4 $\pm$ 0.6 %
-64 91.3 $\pm$ 1.1 % 91.8 $\pm$ 1.3 % 89.6 $\pm$ 1.1 % 89.4 $\pm$ 1.8 %
-32 88.7 $\pm$ 2.2 % 90.2 $\pm$ 1.3 % 86.8 $\pm$ 1.4 % 88.2 $\pm$ 1.2 %
-16 78.1 $\pm$ 1.0 % 86.9 $\pm$ 1.2 % 86.6 $\pm$ 0.7 % 87.1 $\pm$ 1.2 %
-8 60.1 $\pm$ 7.4 % 75.3 $\pm$ 1.6 % 77.3 $\pm$ 2.2 % 80.3 $\pm$ 1.3 %
-(b) $\bm{A}_{sd}$ trainable.
-$h$ $n_{d}=10$ $n_{d}=50$ $n_{d}=100$
-512 92.7 $\pm$ 0.9 % 90.8 $\pm$ 1.5 % 90.4 $\pm$ 1.0 %
-256 92.1 $\pm$ 0.5 % 90.6 $\pm$ 1.3 % 89.9 $\pm$ 0.6 %
-128 92.1 $\pm$ 1.3 % 89.8 $\pm$ 1.2 % 89.4 $\pm$ 0.6 %
-64 91.8 $\pm$ 1.3 % 89.6 $\pm$ 1.1 % 89.4 $\pm$ 1.8 %
-32 90.2 $\pm$ 1.3 % 86.8 $\pm$ 1.4 % 88.2 $\pm$ 1.2 %
-16 86.2 $\pm$ 1.0 % 84.1 $\pm$ 1.2 % 83.0 $\pm$ 1.0 %
-8 72.6 $\pm$ 2.5 % 70.1 $\pm$ 2.6 % 73.8 $\pm$ 1.9 %
-VI-F Trainable vs non-trainable delay parameters
-In this section, we compare the results obtained with trainable and non-trainable $\bm{A}_{sd}$ matrices.
-Comparing Tables III(a) and III(b), we observe that the results are generally within the standard deviation of each other, except for the LIF case, where the trainable $\bm{A}_{sd}$ provides a clear performance benefit.
-Similarly, Tables V(a) and V(b) show no meaningful performance difference between the two configurations.
-Tables IV(a) and IV(b) indicate that the additional training time required when $\bm{A}_{sd}$ is trainable is negligible <1<1 minute, which further supports the discussion in Section IV-D1 regarding the small computational complexity overhead when using our proposed delays.
-VII Discussions
-The results suggest that the LIF and adLIF neurons may not be fully capable of exploiting the introduced delay mechanism.
-Their limited state dimensionality and small number of free parameters likely restrict the model’s ability to fully utilize temporal depth introduced by the delay mechanism.
-More complex neuron models with richer internal dynamics might benefit more from the proposed delay formulation.
-We did not perform any dedicated hyperparameter optimization (HPO) in this study.
-Conducting a thorough HPO for each neuron type and delay configuration could further improve performance and potentially alter observations.
-From a neuromorphic computing perspective, an important open question is which of the proposed delay neuron models can be implemented efficiently on hardware.
-Implementing delays on hardware has been explored [10181778, DelGrad, meijer2025efficientsynapticdelayimplementation, DenRAM], programmable delays can be implemented on various hardware including Intel Loihi [8259423], IBM TrueNorth [7229264], SpiNNaker [6750072], and SENeCA [9870025].
-Whether similar implementations can be performed for the proposed models is an open question.
-VIII Conclusion
-In this work, we proposed a general framework for modeling delays within spiking neuron dynamics by introducing additional state variables that encodes delayed information.
-This formulation enables neurons to access a finite temporal history.
-The framework is general and can be easily integrated into various spiking neuron models.
-We analyzed different aspects of this formulation, including the order of delays, delay parameters and their optimization (trainable or fixed), and the trade-off between delay order and network size.
-Experimental results show that the proposed delay mechanism achieves performance comparable to that of existing delay-based approaches while remaining computationally efficient.
-The proposed delay mechanism introduces either no additional trainable parameters or only a small number relative to the total network size.
-Furthermore, we observe that incorporating delays significantly improves performance in smaller networks, where the delays can compensate for the restricted inherent capacity of the model due to the limited number of neurons.
-IX Appendix
-IX-A Network and Training further details
-Our code is based on the Sparch implementation [bittar2022surrogate].
-During the training process,neuron parameters are clipped to remain within these boundaries as in [deckers2024co] i.e.
+The inputs $i_{s}[t]$ and $i_{d}[t]$ are as defined for RLIF in Section 0.A.1.1 with a correspondence between LIF and adLIF, and RLIF and Recurrent adLIF (RadLIF).
+0.A.2 Linear SSM with Time Shift State Transition Matrix Demonstration
+For the SSM state variable $\bm{v}_{d}$ defined with Time Shift State Transition Matrix and initial state $\bm{v}_{d}[0]=\bm{0}^{n_{d}\times 1}$ we have:
+$\bm{v}_{d}[1]$ $=[i_{d}[1],\,0,\,0,\,...,\,0]^{T}$
+$\bm{v}_{d}[2]$ $=[i_{d}[2],\,a_{1}i_{d}[1],\,0,\,...,\,0]^{T}$
+$\bm{v}_{d}[3]$ $=[i_{d}[3],\,a_{1}i_{d}[2],\,a_{2}a_{1}i_{d}[1],\,...,\,0]^{T}$
+$\,\,\,\vdots$
+$\bm{v}_{d}[t]$ $=[i_{d}[t],\,a_{1}i_{d}[t-1],\,a_{2}a_{1}i_{d}[t-2],\,$...,\,
+$\hskip 85.35826pt(\prod^{n-1}_{k=1}a_{k})i_{d}[t-n_{d}+1]]^{T}.$
+0.A.3 Delay parameters: temporal convolution representation
+Under our parameter choices, the delay state $\bm{v}_{d}[t]$ stores the past $n_{d}$ inputs $i_{d}[t]$.
+The term $\bm{A}_{sd}\bm{v}_{d}[t]$ therefore represents a weighted combination of these past inputs.
+Illustration shown in Figure 4.
+Figure 4: Temporal convolution representation of the delay component.
+Under our parameter choices, the delay variable $\bm{v}_{d}[t]$ stores the past $n_{d}$ inputs $i_{d}[t]$, hence the term $\bm{A}_{sd}\bm{v}_{d}[t]$ corresponds to a weighted combination of them.
+0.A.4 Network and Training further details
+Our code is based on the Sparch implementation [1].
+During the training process,neuron parameters are clipped to remain within these boundaries as in [7] i.e.
 $\alpha\in[0.36,0.96],\beta\in[0.96,0.99],a\in[0,1]$ and $b\in[0,2]$.
-We use hyperparameters based on previous works [hammouamri2023learning, deckers2024co, schöne2024scalableeventbyeventprocessingneuromorphic], in particular base LR $10^{-2}$, WD $10^{-5}$, dropout rate =0.4=0.4, batch size =128=128, and number of epochs =50=50.
+For training we use AdamW optimizer [23] and cosine learning rate scheduler.
+Batch normalization is used following [18].
+We use hyperparameters based on previous works [17, 7, 29], in particular base LR $10^{-2}$, WD $10^{-5}$, dropout rate =0.4=0.4, batch size =128=128, and number of epochs =50=50.
 No hyperparameter optimization has been performed.
-IX-B Dataset further details
-For the SHD dataset, similar to other works [hammouamri2023learning] [deckers2024co] we used spatio-temporal bins to reduce the input dimensions.
+0.A.5 Dataset further details
+For the SHD dataset, similar to other works [17] [7] we used spatio-temporal bins to reduce the input dimensions.
 The input channels were reduced from $700$ to $140$ by binning every $5$ consecutive channels.
-The dataset was put in frames using the To-Frame [tonic] function with time-window=10000=10000, i.e.
+The dataset was put in frames using the To-Frame [22] function with time-window=10000=10000, i.e.
 accumulation of events with time-windows of 1010ms.
 The framed data set was used to load batches during both training and testing.
-During training, we use the data augmentation techniques TimeNeurons_mask_aug and CutMix following previous work [deckers2024co] [hammouamri2023learning].
+During training, we use the data augmentation techniques TimeNeurons_mask_aug and CutMix following previous work [7] [17].